### v2

# Demystifying Spatial Confounding

Abstract
Spatial confounding is a fundamental issue in spatial regression models which arises because spatial random effects, included to approximate unmeasured spatial variation, are typically not independent of covariates in the model. This can lead to significant bias in covariate effect estimates. The problem is complex and has been the topic of extensive research with sometimes puzzling and seemingly contradictory results. Here, we develop a broad theoretical framework that brings mathematical clarity to the mechanisms of spatial confounding, providing explicit analytical expressions for the resulting bias. We see that the problem is directly linked to spatial smoothing and identify exactly how the size and occurrence of bias relate to the features of the spatial model as well as the underlying confounding scenario. Using our results, we can explain subtle and counter-intuitive behaviours. Finally, we propose a general approach for dealing with spatial confounding bias in practice, applicable for any spatial model specification. When a covariate has non-spatial information, we show that a general form of the so-called spatial+ method can be used to eliminate bias. When no such information is present, the situation is more challenging but, under the assumption of unconfounded high frequencies, we develop a procedure in which multiple capped versions of spatial+ are applied to assess the bias in this case. We illustrate our approach with an application to air temperature in Germany.
Keywords: Confounding bias, spatial regression, spatial random effects, smoothing, spatial+.
1 Introduction
In spatial statistics, regression models usually take the form
$y(s)=\beta x(s)+\gamma(s)+\epsilon(s)$ (1)
where $y(s)$ denotes the response variable of interest, $x(s)$ a covariate with regression coefficient $\beta$, $\gamma(s)$ a spatially correlated random effect and $\epsilon(s)$ independent identically distributed error terms, all evaluated at locations $s$ within some spatial domain. We consider the case of only one covariate but the model could be expanded to include more. The spatial effect $\gamma(s)$ is included to approximate unmeasured spatial variation in the response which typically arises due to unobserved spatial variables. Specific instances of the model are obtained by specifying the type of spatial information (e.g. discrete regional vs. continuous coordinate-based) and the representation of the spatial effect $\gamma(s)$ (e.g. Gaussian random fields, Markov random fields, thin plate splines, etc.).
Although not always explicitly stated, the regression (1) assumes that the covariate $x(s)$ is given and fixed. This implies independence between the covariate and the spatial effect, which is unlikely to hold for observational data where covariates are realisations of potentially spatially dependent random variables. If indeed this assumption is violated, it leads to a phenomenon known as spatial confounding where the estimate of the covariate effect $\beta$ becomes biased even when the model is correctly specified. Thus, understanding spatial confounding is fundamental to reliable covariate effect estimation in spatial statistics, and the problem has received considerable attention in the literature since the seminal work of Clayton et al., (1993) and, later, Hodges and Reich, (2010). Recently, there has been a surge in interest, not least because the established understanding and methods most commonly used for dealing with the problem have been shown to be problematic.
Broadly, research into spatial confounding can be structured along the following lines: (i) Investigations and suggestions for potential solutions under specific instances of the model (1) such as a choice of discrete or continuous spatial dependence and/or a particular representation of the spatial effect (Dupont et al.,, 2022; Guan et al.,, 2023; Marques et al.,, 2022; Urdangarin Iztueta et al.,, 2022). (ii) Identifying factors that influence the size or occurrence of bias, in particular, the importance of the spatial scales of both measured and unmeasured variables (Paciorek,, 2010; Page et al.,, 2017; Keller and Szpiro,, 2020; Nobre et al.,, 2021). (iii) Research based on a (sometimes implicit) preference for the estimate of $\beta$ in a non-spatial analysis (Reich et al.,, 2006; Hanks et al.,, 2015; Hughes and Haran,, 2013; Hefley et al.,, 2017; Adin et al.,, 2023; Briz-Redón,, 2023). This approach, usually involving a method known as restricted spatial regression, dominated the literature as well as practical implementations for many years, but has since been shown to be problematic for several reasons (Khan and Calder,, 2022; Zimmerman and Ver Hoef,, 2022). (iv) Research based on causal inference approaches (Thaden and Kneib,, 2018; Papadogeorgou et al.,, 2019; Schnell and Papadogeorgou,, 2020), an overview of which is in Reich et al., (2021).
Although this work has shed more light on spatial confounding, results appear at times to be counter-intuitive or mutually conflicting. This is in part because the complexity of spatial models means that conclusions are often based on simulation studies, potentially sensitive to the specific simulation setup. Moreover, the research direction (iii) above has lead to a focus on the difference between the estimates in the spatial and non-spatial analyses for identifying confounding bias. However, as our results confirm, this is not directly related to the occurrence or otherwise of bias. There have also been attempts to prove that spatial confounding bias in models of the form (1) is always negligible so that the issue does not exist in practice (Khan and Berrett,, 2023), but we will see that bias from spatial confounding can in fact become arbitrarily large.
We develop a theoretical framework that brings mathematical clarity to the mechanisms of spatial confounding and the methods used for dealing with the resulting bias. We see that it is a problem for spatial models in general, directly linked to spatial smoothing (in line with work on bias in linear mixed models, Schnell and Bose,, 2019). Central to our findings is the analytical expression for the bias derived in Proposition 1 which shows that, in the metric defined by the precision matrix of the chosen spatial analysis model, the bias is the correlation between the covariate and the confounder times the size of the confounder relative to the size of the covariate. Not only is this an intuitive result, but by using an eigendecomposition of the precision matrix, we can see how the bias depends on the underlying confounding scenario and the covariance structure and parameters of the analysis model. We also see explicitly the role of spatial frequencies, defined directly in terms of the model. While the expression for the bias is relatively simple, the dependence on individual components and parameters can be complex, explaining the subtle and at times counter-intuitive behaviours. We illustrate our findings in a simulation study of several different scenarios.
Based on our results, we develop a general framework for assessing and adjusting spatial confounding bias in practice, applicable under any specification of the spatial effects in (1). The proposed method depends on whether or not the covariate has non-spatial information. When non-spatial information is present, we show that a general form of spatial+ (Dupont et al.,, 2022) can be used to eliminate bias. When no such information is present, the situation is more challenging as further assumptions for identifiability are needed. Under the assumption of unconfounded high frequencies (as proposed in Guan et al., (2023)), we develop a method that utilises the spatial frequency behaviours of the covariate to assess the bias in this case. Specifically, we propose a procedure in which multiple capped versions of spatial+ are applied.
We illustrate our approach with an application to monthly mean air temperature across Germany during the year 2010 with rainfall as covariate. Both variables display varying spatial frequency behaviours for different months of the year. We show how our methods can be applied to assess whether the effect estimate in the spatial model is biased and, if so, determine the estimated bias.
2 Definitions and preliminary results
2.1 Model specification
Throughout this paper we assume that the true data generating process is given by
$\mathbf{y}=\beta\mathbf{x}+\mathbf{z}+\bm{\epsilon},\quad\bm{\epsilon}\sim N(\mathbf{0},\sigma^{2}\mathbf{I})$ (2)
where $\mathbf{y}=(y(s_{1}),\ldots,y(s_{n}))^{T}$, $\mathbf{x}=(x(s_{1}),\ldots,x(s_{n}))^{T}$ and $\bm{\epsilon}=(\epsilon(s_{1}),\ldots,\epsilon(s_{n}))^{T}$ are the response, covariate and error term at locations $s_{1},\ldots,s_{n}$ and $\mathbf{z}=(z(s_{1}),\ldots,z(s_{n}))^{T}$ the true confounding variable at these locations. Since the confounder is assumed to be unknown, in the spatial analysis model, chosen by the researcher analysing the data, we assume that $\mathbf{z}(s)$ is approximated by a spatial effect $\gamma(s)$ as in (1). This implies an analysis model of the form
$\mathbf{y}=\beta\mathbf{x}+\mathbf{B}_{\mathrm{sp}}\bm{\beta}_{\mathrm{sp}}+\bm{\epsilon},\quad\bm{\beta}_{\mathrm{sp}}\sim N(\mathbf{0},\lambda^{-1}\mathbf{S}^{-}),\text{ }\bm{\epsilon}\sim N(\mathbf{0},\sigma^{2}\mathbf{I})$ (3)
where $\mathbf{B}_{\mathrm{sp}}$ is an $n\times p$ design matrix for the basis representing the spatial effect $\gamma(s)$, and $\bm{\beta}_{\mathrm{sp}}$ the coefficients for the spatial effect in this basis. The matrix $\mathbf{S}$ (with pseudo-inverse $\mathbf{S}^{-}$) is a $p\times p$ penalty matrix which, together with the spatial basis in the design matrix $\mathbf{B}_{\mathrm{sp}}$, defines the structure of the spatial effect, and $\lambda>0$ is a smoothing parameter which controls the overall level of smoothing of the spatial effect. If the model matrix is identifiable (a discussion of this is in Section 5), we can also allow the value $\lambda=0$, in which case the spatial random effect becomes unsmoothed additional fixed effect terms in the model. We assume that $\mathbf{S}$ is symmetric positive semi-definite, i.e. it has non-negative eigenvalues $\alpha_{1}\leq\cdots\leq\alpha_{p}$. The null space of $\mathbf{S}$, which could be non-trivial, corresponds to unsmoothed spatial vectors. The intercept in model (3) is not included as a separate fixed effect, as it can be modelled as part of the spatial effect by including the vector $(1,\ldots,1)^{T}$ as an unpenalised spatial basis vector in $\mathbf{B}_{\mathrm{sp}}$.
For comparison, we also consider the non-spatial analysis model
$\mathbf{y}=\beta_{0}+\beta\mathbf{x}+\bm{\epsilon},\quad\bm{\epsilon}\sim N(\mathbf{0},\sigma^{2}\mathbf{I})$ (4)
where $\beta_{0}\in\mathbb{R}$ is an unknown intercept.
Owing to the generality of our setup, all the commonly used spatial models fit into this generic framework. For example, for (i) Gaussian random fields based on coordinate information, $\mathbf{B}_{\mathrm{sp}}$ is an incidence matrix linking observations to observation locations and $\mathbf{S}$ is the inverse covariance matrix implied by the covariance function (ii) (intrinsic) conditional auto-regressive and Gaussian Markov random field specifications for regional data, $\mathbf{B}_{\mathrm{sp}}$ is an indicator matrix for the regions and $\mathbf{S}$ encodes spatial proximity with an adjacency matrix or graph Laplacian, and (ii) tensor product P-splines and thin plate splines, $\mathbf{B}_{\mathrm{sp}}$ contains basis function evaluations and $\mathbf{S}$ the corresponding smoothness penalty. We can also take a Bayesian perspective if we have a multivariate Gaussian prior on $\bm{\beta}_{\mathrm{sp}}$ with precision $\lambda\mathbf{S}$, see Fahrmeir and Kneib, (2011) for details on both the various options for spatial smoothing and the Bayesian perspective on spatial regression.
2.2 The spatial precision matrix $\bm{\mathbf{\Sigma}}^{-1}$
Central to our further analyses are the properties of the precision matrix $\bm{\mathbf{\Sigma}}^{-1}$ of the spatial analysis model (3), derived from the covariance structure $\bm{\mathbf{\Sigma}}=\sigma^{2}\mathbf{I}+\lambda^{-1}\mathbf{B}_{\mathrm{sp}}\mathbf{S}^{-}\mathbf{B}_{\mathrm{sp}}^{T}$. Details of our derivations are included in Web Appendix 1.1.
Lemma 1.
Let $\alpha_{1}\leq\cdots\leq\alpha_{p}$ be the eigenvalues of the penalty matrix $\mathbf{S}$ and $\lambda>0$ the smoothing parameter. Then the eigenvalues of the precision matrix $\bm{\mathbf{\Sigma}}^{-1}$ are given by $\{\sigma^{-2},\sigma^{-2}w_{1},\ldots,\sigma^{-2}w_{p}\}$ where $w_{i}=\lambda\alpha_{i}/(\sigma^{-2}+\lambda\alpha_{i})\textrm{ for }i=1,\ldots,p.$
Lemma 1 shows that there exists an $n\times n$-matrix $\mathbf{U}$ such that $\mathbf{U}^{T}\bm{\mathbf{\Sigma}}^{-1}\mathbf{U}=\sigma^{-2}\textrm{diag}(1,\ldots,1,w_{1},\allowbreak\ldots,w_{p})$ where $w_{i}=\lambda\alpha_{i}/(\sigma^{-2}+\lambda\alpha_{i})$ for $i=1,\dots,p$. The columns of $\mathbf{U}$ are an orthonormal basis of eigenvectors. The first $n-p$ eigenvectors $\mathbf{U}_{\mathrm{ns}}$ span the non-spatial part of $\mathbb{R}^{n}$ which is unaffected by smoothing. The remaining $p$ eigenvectors $\mathbf{U}_{\mathrm{sp}}$ represent different types of spatial behaviour, where the $i$th spatial eigenvector has eigenvalue $\sigma^{-2}w_{i}$ with $w_{i}$ a weight between 00 and $1$, directly linked to the smoothing penalty $\alpha_{i}$. Since $w_{1}\leq\cdots\leq w_{p}$, it is the spatial vectors with higher weights that correspond to behaviours that are penalised more, i.e. that are considered “less spatial” according to the analysing model. Thus, we will refer to behaviour given by a spatial eigenvector with weight $w_{i}$ close to 1 as being “high frequency” and eigenvectors with low weights $w_{i}$ as “low frequency”. In this way, what is considered high/low frequency spatial behaviour is always well-defined and directly linked to the spatial structure chosen for the spatial analysis model. Since the precision matrix $\bm{\mathbf{\Sigma}}^{-1}$ is symmetric positive definite it defines an inner product on $\mathbb{R}^{n}$. For the rest of this paper, if $\mathbf{M}$ is a symmetric positive definite $n\times n$ matrix, we will use the notation $\langle\cdot,\cdot\rangle_{\mathbf{M}}$ to denote the inner product defined by $\mathbf{M}$ and $\|\cdot\|_{\mathbf{M}}$ the corresponding norm. That is, $\langle\mathbf{v},\mathbf{v}^{\prime}\rangle_{\mathbf{M}}=\mathbf{v}^{T}\mathbf{M}\mathbf{v}^{\prime}$ and $\|\mathbf{v}\|_{\mathbf{M}}=\sqrt{\langle\mathbf{v},\mathbf{v}\rangle_{\mathbf{M}}}$ for all $\mathbf{v},\mathbf{v}^{\prime}\in\mathbb{R}^{n}$.
3 Characterising the impact of spatial confounding
In this section we present the central results of the paper, namely, Proposition 1 and Corollary 1, which provide explicit expressions for the bias of the covariate effect estimate in the analysis model (3) arising from spatial confounding.
We assume throughout that the true data generating process is that of model (2) with true covariate effect $\beta$ and true spatial confounder $\mathbf{z}$, where the latter is assumed to be generated from the spatial structure of model (3). Thus, our results do not rely on model mis-specification; the bias occurs even when the analysis model is perfectly able to reproduce the confounder. Further bias could arise when the analysis model is mis-specified, but this is not our focus here.
3.1 Bias in the spatial model
Proposition 1.
The bias of the estimated covariate effect $\widehat{\beta}$ in model (3) is given by
$\operatorname{E}(\widehat{\beta})-\beta=\frac{\langle\mathbf{x},\mathbf{z}\rangle_{\bm{\mathbf{\Sigma}}^{-1}}}{\langle\mathbf{x},\mathbf{x}\rangle_{\bm{\mathbf{\Sigma}}^{-1}}}=\frac{\langle\mathbf{x},\mathbf{z}\rangle_{\bm{\mathbf{\Sigma}}^{-1}}}{\|\mathbf{x}\|_{\bm{\mathbf{\Sigma}}^{-1}}\|\mathbf{z}\|_{\bm{\mathbf{\Sigma}}^{-1}}}\frac{\|\mathbf{z}\|_{\bm{\mathbf{\Sigma}}^{-1}}}{\|\mathbf{x}\|_{\bm{\mathbf{\Sigma}}^{-1}}}.$
Thus, the bias has the intuitive interpretation as the correlation between $\mathbf{x}$ and $\mathbf{z}$ times the size of $\mathbf{z}$ relative to the size of $\mathbf{x}$ in the metric defined by $\bm{\mathbf{\Sigma}}^{-1}$. That is, in this metric, the bias gets large exactly when the covariate and the confounder are very correlated and the confounder is relatively large compared to the covariate. The sign of the bias is the same as the sign of the correlation, i.e. the sign of $\langle\mathbf{x},\mathbf{z}\rangle_{\bm{\mathbf{\Sigma}}^{-1}}$. The proof of the proposition is in the appendix.
Combining Proposition 1 with the results of Section 2 we get a better understanding of how the bias behaves. Let $\bm{\xi}^{x}=(\bm{\xi}^{xT}_{\mathrm{ns}},\bm{\xi}^{xT}_{\mathrm{sp}})^{T}$ denote the coordinates of $\mathbf{x}$ in the eigenbasis $\mathbf{U}$ such that $\mathbf{x}=\mathbf{U}\bm{\xi}^{x}$, i.e. a non-zero entry in $\bm{\xi}^{x}$ means that the corresponding basis vector is included in $\mathbf{x}$. Similarly, let $\mathbf{z}=\mathbf{U}\bm{\xi}^{z}$. Since $\mathbf{z}$ is spatial, $\bm{\xi}^{z}_{\mathrm{ns}}=\mathbf{0}$. It is now straightforward to prove the following result (see the Web Appendix for the proof).
Corollary 1.
Let $\mathbf{U}$ be the orthonormal eigenbasis which diagonalises $\bm{\mathbf{\Sigma}}^{-1}$, and $\bm{\xi}^{x}=(\bm{\xi}^{xT}_{\mathrm{ns}},\bm{\xi}^{xT}_{\mathrm{sp}})^{T}$ and $\bm{\xi}^{z}=(\mathbf{0}^{T},\bm{\xi}^{zT}_{\mathrm{sp}})^{T}$ the coordinates of $\mathbf{x}$ and $\mathbf{z}$ in this basis. The bias of the estimated covariate effect $\widehat{\beta}$ in model (3) is given by
$\operatorname{E}(\widehat{\beta})-\beta=\frac{\sum_{i=1}^{p}\xi_{\mathrm{sp},i}^{x}\xi_{\mathrm{sp},i}^{z}w_{i}}{\sum_{i=1}^{n-p}(\xi_{\mathrm{ns},i}^{x})^{2}+\sum_{i=1}^{p}(\xi_{\mathrm{sp},i}^{x})^{2}w_{i}},$
where $w_{i}=\lambda\alpha_{i}/(\sigma^{-2}+\lambda\alpha_{i})$ for $i=1,\ldots,p$.
The expression in Corollary 1 (consistent with (10) of Schnell and Bose, (2019)) shows that bias in the spatial model is directly linked to spatial smoothing, as without smoothing (i.e. if $\alpha_{i}=0$ for all $i$ or $\lambda=0$), the bias would not arise. In other words, adding spatial random effects to the model (4) adjusts for unmeasured spatial confounders thereby, in principle, eliminating bias, however, bias is reintroduced as a result of spatial smoothing. For any given $\lambda>0$, i.e. a fixed overall level of smoothing, we see that bias occurs whenever $\mathbf{x}$ and $\mathbf{z}$ share a spatial component for which $w_{i}\neq 0$. More specifically, any confounded component in $\mathbf{x}$, i.e. with $\xi_{\mathrm{sp},i}^{x}\xi_{\mathrm{sp},i}^{z}\neq 0$, contributes to the numerator of the bias. At the same time, all components in $\mathbf{x}$ increase the denominator, thereby reducing the overall size of the bias.
The expression for the bias is not symmetric in $\mathbf{x}$ and $\mathbf{z}$. In particular, it is not affected by components in $\mathbf{z}$, i.e. with $\xi_{\mathrm{sp},i}^{z}\neq 0$, for which the corresponding component in $\mathbf{x}$ is 0. Thus, any unmeasured spatial variation in the response variable that is independent of the covariate can be ignored for the purposes of the bias. However, such components may affect the overall fit of the model and the estimates of the parameters $\lambda$ and $\sigma^{2}$ which affect the weights $w_{i}$.
3.2 The role of spatial frequencies
Due to the differences in the weights $w_{i}$, different spatial frequencies affect the bias differently. Looking at the numerator in Corollary 1, the larger the weight, the larger the contribution to the bias. So confounding at high frequencies (where weights are close to $1$) induce larger bias, while confounding at low frequencies may give negligible contribution to the bias. Confounding at unpenalised spatial frequencies does not lead to bias (as they have weights 00), and non-spatial components in $\mathbf{x}$ do not contribute to the numerator either. For all confounded frequencies in $\mathbf{x}$ with $w_{i}\neq 0$, the contribution to the bias is proportional to the size $\xi_{\mathrm{sp},i}^{z}$ of the confounder at that frequency. Contributions to the bias from different eigenvectors can partially cancel each other out as they can have different signs.
From the denominator we see that any unconfounded components of $\mathbf{x}$, that is, either non-spatial components $\xi_{\mathrm{ns},i}^{x}\neq 0$ or spatial components with $\xi_{\mathrm{sp},i}^{x}\neq 0$ and $\xi_{\mathrm{sp},i}^{z}=0$, will reduce the size of the bias, as they increase the denominator without affecting the numerator. This reduction will be larger when the weight $w_{i}$ is large, i.e. when the spatial frequency is high, and may be negligible or even 00 for low frequencies. Unconfounded high frequency spatial components of $\mathbf{x}$ have a similar effect on the bias as non-spatial components (in line with the intuition that high frequencies are “less spatial”), i.e. they contribute nothing to the numerator and are included with weight close to $1$ in the denominator.
3.3 Comparison to the non-spatial model
Irrespective of the size of bias, the non-spatial model (4) is not an appropriate analysis model for the data generation process (2), as the residual spatial variation in the response data violates the assumptions of the model. Indeed, this is the reason for including spatial random effects. Nevertheless, we consider how the bias in model (4) compares to the spatial model (3) as this has been a focus of spatial confounding research in the past. The following corollary is proved in the appendix.
Corollary 2.
Let $\mathbf{U}$ be the orthonormal eigenbasis which diagonalises $\bm{\mathbf{\Sigma}}^{-1}$, and $\bm{\xi}^{x}=(\bm{\xi}^{xT}_{\mathrm{ns}},\bm{\xi}^{xT}_{\mathrm{sp}})^{T}$ and $\bm{\xi}^{z}=(\mathbf{0}^{T},\bm{\xi}^{zT}_{\mathrm{sp}})^{T}$ the coordinates of $\mathbf{x}$ and $\mathbf{z}$ in this basis. The bias of the estimated covariate effect $\widehat{\beta}_{\mathrm{ns}}$ in model (4) is given by
$\operatorname{E}(\widehat{\beta}_{\mathrm{ns}})-\beta=\frac{\langle\mathbf{x},\mathbf{z}\rangle}{\langle\mathbf{x},\mathbf{x}\rangle}=\frac{\langle\mathbf{x},\mathbf{z}\rangle}{\|\mathbf{x}\|\|\mathbf{z}\|}\frac{\|\mathbf{z}\|}{\|\mathbf{x}\|}\\ =\frac{\sum_{i=1}^{p}\xi_{\mathrm{sp},i}^{x}\xi_{\mathrm{sp},i}^{z}}{\sum_{i=1}^{n-p}(\xi_{\mathrm{ns},i}^{x})^{2}+\sum_{i=1}^{p}(\xi_{\mathrm{sp},i}^{x})^{2}}.$
Thus, the bias in the non-spatial model differs to that of the spatial model only in that all spatial frequencies have weights $1$ rather than $w_{i}$. Broadly speaking, bias in the non-spatial model reflects the overall correlation between $\mathbf{x}$ and $\mathbf{z}$, whereas bias in the spatial model depends mostly on the correlation at high frequencies. However, as the weights affect both the numerator and denominator of the bias, in general, the relationship between the two models is not straightforward. A detailed computation of the difference and how it is affected by different types of behaviours is given in Web Appendix 1.4. Although the bias tends to be larger in the non-spatial model than in the spatial model, there are scenarios where it is the other way around.
In the past, the estimate in model (4) was often mistakenly assumed to be unbiased and, therefore, a large difference between this and the estimate in (3) was seen as a sign of confounding bias in the spatial model (and, conversely, a small difference a sign of no confounding). However, Corollary 2 shows that a difference between these estimates tends to arise when the covariate is dominated by low frequencies (as these have weights most different from 1) and is not directly related to the existence or otherwise of confounding bias.
3.4 Dependence on the smoothing parameter
As bias in the spatial model is caused by smoothing, intuitively, we would expect larger values of the parameter $\lambda$ to lead to more bias when all other components are held fixed. However, as $\lambda$ affects both the numerator and the denominator of the expression in Corollary 1, there can be situations where this is not the case. In Web Appendix 1.5 we provide a detailed analysis which shows that, although the behaviour of the bias varies, when there is sufficient non-spatial or unconfounded high frequency components in $\mathbf{x}$, the size of the bias broadly increases with $\lambda$ (in line with intuition). The following corollary is also proved in the Web Appendix.
Corollary 3.
Let $\mathbf{U}$ be the orthonormal eigenbasis which diagonalises $\bm{\mathbf{\Sigma}}^{-1}$, and $\bm{\xi}^{x}=(\bm{\xi}^{xT}_{\mathrm{ns}},\bm{\xi}^{xT}_{\mathrm{sp}})^{T}$ and $\bm{\xi}^{z}=(\mathbf{0}^{T},\bm{\xi}^{zT}_{\mathrm{sp}})^{T}$ the coordinates of $\mathbf{x}$ and $\mathbf{z}$ in this basis. The bias of the estimated covariate effect $\widehat{\beta}$ in model (3) has the following limiting behaviour:
$\lim_{\lambda\rightarrow 0}\operatorname{E}(\widehat{\beta})-\beta$ $=$ $\left\{\begin{array}[]{cr}0&\text{if }\bm{\xi}_{\mathrm{ns}}^{x}\neq\mathbf{0}\\ \frac{\sum_{i=1}^{p}\xi_{\mathrm{sp},i}^{x}\xi_{\mathrm{sp},i}^{z}\alpha_{i}}{\sum_{i=1}^{p}(\xi_{\mathrm{sp},i}^{x})^{2}\alpha_{i}}&\text{otherwise}\end{array}\right.$
$\lim_{\lambda\rightarrow\infty}\operatorname{E}(\widehat{\beta})-\beta$ $=$ $\frac{\sum_{\{i|\alpha_{i}\neq 0\}}\xi_{\mathrm{sp},i}^{x}\xi_{\mathrm{sp},i}^{z}}{\sum_{i=1}^{n-p}(\xi_{\mathrm{ns},i}^{x})^{2}+\sum_{\{i|\alpha_{i}\neq 0\}}(\xi_{\mathrm{sp},i}^{x})^{2}}.$
Thus, in the limit $\lambda\rightarrow\infty$, i.e. an entirely smoothed spatial effect, the bias in the spatial model agrees with that of the non-spatial model, except that any contributions from unpenalised spatial components of $\mathbf{x}$ are 00.
In the limit $\lambda\rightarrow 0$, i.e. without smoothing, the behaviour of the bias depends on whether or not the covariate has non-spatial information. If non-spatial information is present, the bias can be eliminated by setting $\lambda=0$, though this is usually undesirable as it is likely lead to an inferior model fit. Similarly, if the proportion of unconfounded high frequency components in $\mathbf{x}$ is large, then the size of the bias reduces as $\lambda\rightarrow 0$ and the limit is close to 00. But in the absence of non-spatial or unconfounded high frequency information, the limit could become large and even exceed the bias of the non-spatial model.
4 Simulation studies
In the following sections we simulate different confounding scenarios to study the behaviour of the bias of the covariate effect estimate $\widehat{\beta}$ in model (3) and $\widehat{\beta}_{\mathrm{ns}}$ in model (4). For each scenario we generate 100 independent replicates of covariate data $\mathbf{x}$ and response data $\mathbf{y}$, observed at $n=1000$ randomly selected locations in the spatial domain $[0,1]\times[0,1]\subset\mathbb{R}^{2}$ such that
$\mathbf{y}=\beta\mathbf{x}+\mathbf{z}+\bm{\epsilon}^{y},\qquad\mathbf{x}=\mathbf{z}^{x}+\bm{\epsilon}^{x}$
where $\beta\in\mathbb{R}$, $\bm{\epsilon}^{y}\sim N(\mathbf{0},\sigma^{2}\mathbf{I})$, $\bm{\epsilon}^{x}\sim N(\mathbf{0},\sigma^{2}_{x}\mathbf{I})$ for $\sigma,\sigma_{x}>0$. In this specification, $\bm{\epsilon}^{x}$ represents the non-spatial component of $\mathbf{x}$, while $\mathbf{z}$ and $\mathbf{z}^{x}$ are constructed as linear combinations of two spatial fields $\mathbf{z}_{\mathrm{sp,l}}$ and $\mathbf{z}_{\mathrm{sp,h}}$, which are low and high frequency, respectively, i.e. $\mathbf{z}=\xi_{\mathrm{sp,l}}^{z}\mathbf{z}_{\mathrm{sp,l}}+\xi_{\mathrm{sp,h}}^{z}\mathbf{z}_{\mathrm{sp,h}}$ and $\mathbf{z}^{x}=\xi_{\mathrm{sp,l}}^{x}\mathbf{z}_{\mathrm{sp,l}}+\xi_{\mathrm{sp,h}}^{x}\mathbf{z}_{\mathrm{sp,h}}$ for $\xi_{\mathrm{sp,l}}^{z},\xi_{\mathrm{sp,h}}^{z},\xi_{\mathrm{sp,l}}^{x},\xi_{\mathrm{sp,h}}^{x}\in\mathbb{R}$. Unless stated otherwise, we let $\beta=0.5$, $\sigma_{x}=\sigma=1$.
For the regressions in this section, we represent the spatial effect as a thin plate regression spline (TPRS), a computationally efficient reduced rank version of a thin plate spline (Wood,, 2003). The data generating process matches the analysing spatial model which is therefore correctly specified. (In Web Appendix 4, we repeat the simulation where the data generating process is a Gaussian process, so in addition to the spatial confounding bias, there may be further mis-specification bias.) In order to generate $\mathbf{z}_{\mathrm{sp,l}}$ and $\mathbf{z}_{\mathrm{sp,h}}$, we start by considering a mean-zero Gaussian process $\bm{\gamma}$ with exponential covariance structure following $C(h)=\exp(-h/\kappa)$ such data $h=\lVert\mathbf{s}-\mathbf{s}^{\prime}\rVert$ for $\mathbf{s},\mathbf{s}^{\prime}\in[0,1]\times[0,1]$, which is assumed for simplicity to have variance of 1. We set $\kappa=0.1$ corresponding to a spatial range of approximately 0.3. We fit thin plate regression splines to the generated $\bm{\gamma}$ with 10 and 800 basis functions, respectively, and use the corresponding fitted values as $\mathbf{z}_{\mathrm{sp,l}}$ and $\mathbf{z}_{\mathrm{sp,h}}$. Thin plate regression spline models are implemented in the R-package mgcv. We use this implementation with generalised cross-validation as the smoothness selection criterion to compare the results of models fitted to simulated data for which we know the true underlying covariate and spatial dependence. We consider 600 basis functions for the thin plate regression splines in the data analysing model.
4.1 Scenario 1: Confounding at high frequencies
From Corollary 1 we expect confounding at high frequencies, i.e. $\xi_{\mathrm{sp,h}}^{x}\xi_{\mathrm{sp,h}}^{z}\neq 0$, to lead to bias in the spatial model. The larger the value of $\xi_{\mathrm{sp,h}}^{x}\xi_{\mathrm{sp,h}}^{z}$, the larger the numerator in the bias. In fact, keeping everything else fixed, the bias increases linearly with $\xi_{\mathrm{sp,h}}^{z}$ and can therefore become arbitrarily large. From Corollary 2 we expect the same behaviour in the non-spatial model, as the bias has the same expression as that of the spatial model, except that each weight $w_{i}$ is replaced by $1$. Thus, we consider $(\xi_{\mathrm{sp,l}}^{x},\xi_{\mathrm{sp,h}}^{x})=(1,1)$, $\xi_{\mathrm{sp,l}}^{z}=0$ and $\xi_{\mathrm{sp,h}}^{z}\in\{0.2,0.5,1\}$.
As expected, Figure 1 (top left) shows significant bias in both models under this scenario, increasing linearly in size with $\xi_{\mathrm{sp,h}}^{z}$. The mean squared error (MSE) of fitted values compared to the true expectation of $\mathbf{y}$ is shown in Web Appendix 3.1 and is always lower for the spatial model, indicating a better fit, which is not surprising as the spatial model is correctly specified, whereas the non-spatial model is mis-specified.
Figure 1: Bias of $\widehat{\beta}$ and $\widehat{\beta}_{\mathrm{ns}}$ (top) in the spatial and non-spatial (ns) models under Scenario 1 (top, left), Scenario 2 (top, right), Scenario 3 (bottom, left), Scenario 4 (bottom, right). The true value of $\beta$ is 0.5.
4.2 Scenario 2: Confounding at low frequencies
We reconsider Scenario 1 but with high and low frequencies swapped, i.e. $(\xi_{\mathrm{sp,l}}^{x},\xi_{\mathrm{sp,h}}^{x})=(1,1)$, $\xi_{\mathrm{sp,h}}^{z}=0$ and $\xi_{\mathrm{sp,l}}^{z}\in\{0.2,0.5,1\}$. Since confounding is now at low frequencies, i.e. $\xi_{\mathrm{sp,l}}^{x}\xi_{\mathrm{sp,l}}^{z}\neq 0$, contributions to the numerator of the bias in the spatial model are multiplied by low weights $w_{i}$, while the denominator of the bias is still relatively large due to the non-spatial information in $\mathbf{x}$. Therefore we expect the bias to be small. This is confirmed in the results shown in Figure 1 (top right) where the bias in the spatial model remains close to zero, even when $\xi_{\mathrm{sp,l}}^{z}$ is increased. In contrast, swapping frequencies makes no difference to the non-spatial model as all frequencies have equal weight and, indeed, the behaviour for this model is the same as in Scenario 1. Web Appendix 3.1 confirms, as before, that the spatial model is superior in terms of fit.
4.3 Scenario 3: Comparison to the non-spatial model
In Section 3.3 we noted that difference or similarity between the estimates in the spatial and non-spatial models is not directly linked to the existence or otherwise of bias in the spatial model. This is illustrated in Scenarios 1 and 2 above, as in both of these scenarios the estimates in the two models differ but, in Scenario 1, the spatial model is biased whereas, in Scenario 2, it is not. To illustrate this point further, here we simulate two scenarios in which the estimates in the two models are similar, but in one scenario both models are biased whereas in the other they are both largely unbiased. Specifically, in order to generate $\mathbf{z}_{\mathrm{sp,h}}$ and $\mathbf{z}_{\mathrm{sp,l}}$, we fit a reparametrised TPRS spatial effect (see Web Appendix 2) to the Gaussian process with $\kappa=0.1$ where the corresponding 75 highest frequencies are used to generate $\mathbf{z}_{\mathrm{sp,h}}$ and the remaining frequencies to generate $\mathbf{z}_{\mathrm{sp,l}}$. We consider: (1) biased scenario with $(\xi_{\mathrm{sp,l}}^{x},\xi_{\mathrm{sp,h}}^{x})=(0,4)$ and $(\xi_{\mathrm{sp,l}}^{z},\xi_{\mathrm{sp,h}}^{z})=(1,1)$; (2) unbiased scenario with $(\xi_{\mathrm{sp,l}}^{x},\xi_{\mathrm{sp,h}}^{x})=(0,0.1)$ and $(\xi_{\mathrm{sp,l}}^{z},\xi_{\mathrm{sp,h}}^{z})=(1,0)$. We use $n$ basis functions in the the data analysis model. In the first scenario, $\mathbf{x}$ has only high frequencies which are all confounded, so the spatial model is biased with weights $w_{i}$ close to $1$, making the bias similar to that of the non-spatial model. In the second scenario, there are no overlapping frequencies between $\mathbf{x}$ and $\mathbf{z}$ so both models are unbiased. The results are shown in Figure 1 (bottom, left).
Section 3.3 also showed that while the non-spatial bias tends to exceed the spatial bias, it could also be the other way around. In both Scenarios 1 and 2, the non-spatial bias always stays above the spatial bias. However, in Web Appendix 3.2, we repeat Scenario 1 with a larger proportion of unconfounded low frequency components in $\mathbf{x}$ so that the correlation between $\mathbf{x}$ and $\mathbf{z}$ is low overall, but high at high frequencies. As expected from our analysis, in this latter scenario, the spatial bias exceeds the non-spatial bias.
4.4 Scenario 4: Dependence on non-spatial information
We now study the effect on the bias of non-spatial information in $\mathbf{x}$ by increasing the parameter $\sigma_{x}$. As the coefficient $\xi_{\mathrm{ns}}^{x}$ appears only in the denominator of the bias in both the spatial and the non-spatial model, we expect bias in both models to reduce. Specifically, we consider $\sigma_{x}\in\{0.3,1,3\}$. Other parameters remain constant at $(\xi_{\mathrm{sp,l}}^{x},\xi_{\mathrm{sp,h}}^{x})=(1,1)$, $(\xi_{\mathrm{sp,l}}^{z},\xi_{\mathrm{sp,h}}^{z})=(0,0.5)$.
Figure 1 (bottom, right) shows that, as expected, the bias decreases for both the non-spatial and spatial model when $\sigma_{x}$ is increased. As before, the MSE of fitted values is always larger for the non-spatial model. In Section 3.3, we saw that non-spatial bias tends to exceed that of the spatial model, but if the covariate has little non-spatial information and confounding takes place at high frequencies, then it may be the other way around. This behaviour is observed in Figure 1 where the non-spatial bias is higher than the spatial, except in the case $\sigma_{x}=0.3$, i.e where $\mathbf{x}$ has the smallest amount of non-spatial information.
4.5 Scenario 5: Dependence on the smoothing parameter
As detailed in Section 3, the expression for the bias in the spatial model has a relatively complex dependency on the smoothing parameter $\lambda$. However, when there are sufficient non-spatial or unconfounded high frequency components in $\mathbf{x}$, we expect the bias to broadly increase from zero and, assuming there are not many unpenalised spatial components in $\mathbf{x}$, eventually approach the bias in the corresponding non-spatial model. Here, we simulate this type of data under three data generation processes, chosen to broadcast a range of behaviours. The simulated data has confounding at high frequencies, unconfoundedness at low frequencies and varying amounts of non-spatial information, specifically, $\sigma_{x}\in\{0.3,1,2\}$. We use 10 basis functions to generate $\mathbf{z}_{\mathrm{sp,l}}$ and 800 basis functions for $\mathbf{z}_{\mathrm{sp,h}}$. Moreover, we let $(\xi_{\mathrm{sp,l}}^{x},\xi_{\mathrm{sp,h}}^{x})=(1,1)$, $(\xi_{\mathrm{sp,l}}^{z},\xi_{\mathrm{sp,h}}^{z})=(0,0.5)$ and fix $\lambda$ at different values. We consider 20 replicates and compute the average spatial and non-spatial bias for those replicates using 800 basis functions in the analysing spatial model.
Figure 2 shows that for $\sigma_{x}=0.3$, the spatial bias increases from 0, crosses that of the non-spatial model and stays above it, until they meet again for large $\lambda$. As expected, increasing non-spatial information in $\mathbf{x}$ decreases bias in general, and for $\sigma_{x}=1$ and $\sigma_{x}=2$, the bias of the spatial model stays below that of the non-spatial model. For each data generating process in Figure 2, we also show the median estimated smoothing parameter $\hat{\lambda}$ using the generalised cross-validation criterion. At the estimated values of $\lambda$ the bias is 0.4037, 0.1804 and 0.0362, from left to right. For $\sigma_{x}=0.3$, the generalised cross-validation estimate is in a relatively flat area of the curve where the spatial bias exceeds the non-spatial. For both of the other values of $\sigma_{x}$, although the spatial bias is bounded by the non-spatial bias, the estimate $\hat{\lambda}$ is at the steep part of the curve, which means that small changes in $\hat{\lambda}$ can lead to large changes in the bias.
Figure 2: Bias of $\widehat{\beta}$ (solid black) and $\widehat{\beta}_{\mathrm{ns}}$ (dashed black) for increasing $\lambda$ and different $\sigma_{x}$. The true value of $\beta$ is 0.5. The magenta dashed line shows the median $\hat{\lambda}$ estimated by generalised cross-validation.
5 Bias adjustment
As the expressions in Section 3 involve the unknown confounder $\mathbf{z}$, in practical applications, we cannot simply compute the bias to assess the impact of confounding. As illustrated in Section 3.3 and the Scenario 3 simulations above, we also cannot use the difference between the estimates in the spatial and non-spatial models as an indicator for the existence or otherwise of bias. Here, we propose a general framework for bias assessment and adjustment based on our results. That is, for a given data set and choice of spatial analysis model, we propose methods for estimating the bias under different assumptions for the data generation scenario. One assumption that is not usually highlighted in the spatial confounding literature is whether the covariate $\mathbf{x}$ has non-spatial information. However, as illustrated in Sections 3 and 4, the bias may exhibit fundamentally different behaviour depending on whether or not this assumption holds. As outlined below, this assumption is also important in the context of bias adjustment. hile the work on spatial scales is well-known, the results may be somewhat counter-intuitive, and there has been a tendency to worry more about confounding at low spatial frequencies despite this being the less problematic scenario. This is perhaps because low frequencies are often seen as more likely to be confounded in practice, with higher frequencies expected to capture the more characteristic behaviour of the variables. Moreover, as seen from our comparison between the spatial vs. the non-spatial model, low frequency confounding can also lead to large differences in the estimates in these two models. Historically, such differences were interpreted as a sign of confounding with the non-spatial estimate mistakenly assumed to be unbiased. Our results confirm that bias in the spatial model is unlikely to be a problem when confounding is at low frequencies, as long as there is (sufficient) non-spatial or unconfounded high frequency information in the covariate.
5.1 Case 1: $\mathbf{x}$ has non-spatial information
If the covariate $\mathbf{x}$ has non-spatial information, i.e. $\bm{\xi}^{x}_{\mathrm{ns}}\neq\mathbf{0}$, it means that, while the locations at which the data are collected induce some spatial dependency structure in the covariate data, there is additional variation that cannot be explained by the spatial information alone. In this case, the model matrix for (3) has full rank and the model is identifiable. Thus, the model has information to distinguish the effect of $\mathbf{x}$ from that of $\mathbf{z}$, but bias in the covariate effect estimate arises as a result of distortion caused by spatial smoothing. If our only objective was to avoid the bias, we could simply remove or reduce the level of smoothing, indeed, Corollary 3 shows that the bias becomes zero as $\lambda\rightarrow 0$. However, undersmoothing is undesirable from the point of view of model fit, as smoothing is applied to reflect the spatial covariance structure and avoid overfitting the spatial effect.
Spatial+ (Dupont et al.,, 2022) is a method for bias adjustment which utilises the fact that the non-spatial component $\mathbf{x}_{\mathrm{ns}}=\mathbf{U}_{\mathrm{ns}}\bm{\xi}^{x}_{\mathrm{ns}}$ of the covariate is unconfounded and can therefore be used to identify the covariate effect. Although originally introduced for models in which the spatial effect is a thin plate spline, the method can be formulated for any spatial analysis model as follows. The model (3) with no covariates is initially fitted to $\mathbf{x}$ to give the decomposition $\mathbf{x}=\mathbf{x}_{\mathrm{sp}}+\mathbf{r}^{x}$ where $\mathbf{x}_{\mathrm{sp}}$ are the fitted values (estimating the spatial part of $\mathbf{x}$) and $\mathbf{r}^{x}$ the residuals. The covariate effect $\beta$ is then estimated by fitting the model (3) but with $\mathbf{x}$ replaced by $\mathbf{r}^{x}$. From Proposition 1, the bias in spatial+ is therefore $\langle\mathbf{r}^{x},\mathbf{z}\rangle_{\bm{\mathbf{\Sigma}}^{-1}}/\langle\mathbf{r}^{x},\mathbf{r}^{x}\rangle_{\bm{\mathbf{\Sigma}}^{-1}}$. As the residuals $\mathbf{r}^{x}$ estimate the non-spatial component $\mathbf{x}_{\mathrm{ns}}$ which is orthogonal to $\mathbf{z}$, this shows why we can expect the bias to be eliminated. Indeed, even though $\mathbf{r}^{x}$ may contain a small proportion of potentially confounded spatial components, its non-spatial part is likely to dominate so that bias is negligible (as seen in the Scenario 4 simulations).
5.2 Case 2: $\mathbf{x}$ is fully spatial
Sometimes $\mathbf{x}$ has no non-spatial components, i.e. $\bm{\xi}^{x}_{\mathrm{ns}}=\mathbf{0}$. For example, $\mathbf{x}$ could be entirely generated from a smooth spatial process, or the dimension $p$ of the spatial effect in (3) could be as large as the sample size $n$ (such as in a regional spatial model with only one data measurement per region). The covariate vector is then contained in the space spanned by the spatial basis vectors and model (3) is unidentifiable. Here, the smoothing penalty not only improves the fit but acts as a regulariser. The model cannot be fitted without smoothing, and the apportionment of the covariate effect between the covariate and spatial terms of the model is directly related to how the smoothing is applied. Our expression for the bias still holds, however, unlike Case 1 above where the non-spatial component $\mathbf{x}_{\mathrm{ns}}$ is known to be unconfounded, we have no information of the confounding scenario and methods like spatial+ may not be effective. The behaviour of the bias will also generally be harder to predict. In particular, even if confounding is at low frequencies (making the numerator of the bias small), the bias could still be significant as the absence of non-spatial components in $\mathbf{x}$ means that the denominator could also potentially be very small, especially when the covariate is dominated by low frequency behaviours.
As shown in Sections 3 and 4, spatial frequencies play a key role for the bias, and techniques such as Fourier and wavelet analysis are useful for understanding the behaviours of the observed variables $\mathbf{y}$ and $\mathbf{x}$ (as suggested, for example, in Keller and Szpiro, (2020)). However, our expressions show that it is the frequencies shared by $\mathbf{x}$ and $\mathbf{z}$ that cause bias, and the data has no information on how such frequencies are split between the two variables. Therefore, without further (untestable) assumptions for identifiability, we cannot estimate the bias. In the case where $\mathbf{x}$ and $\mathbf{z}$ are realisations of two stationary Gaussian spatial processes, Guan et al., (2023) use a Fourier transform to describe spatial frequencies and show that an assumption of unconfoundedness at high frequencies can be used to obtain identifiability. As noted in Guan et al., (2023), this assumption seems natural as the highest frequency spatial variation in a variable is more likely to be characteristic to that particular variable and therefore not confounded with other spatial variables. Corollary 1 shows that this idea extends to our framework in which spatial frequencies are defined directly in terms of the spatial analysis model (3) as described in Section 2.2. Unconfoundedness at the highest frequencies then corresponds to the assumption that $\xi^{z}_{\mathrm{sp},i}=0$ for frequencies above a certain threshold. Thus, under this assumption, the frequency components of $\mathbf{x}$ above the threshold can be used to obtain identifiability.
5.3 Capped spatial+
Based on our analysis in Section 5.2, we present an adapted version of spatial+ which can be used to assess the bias in the more challenging Case 2 scenario. As previously noted, unconfounded high frequencies in $\mathbf{x}$ have a similar effect on the bias as non-spatial information. Therefore, assuming the $k$ highest frequencies are unconfounded, a natural way to adapt spatial+ is to (i) decompose the covariate as $\mathbf{x}=\mathbf{x}_{\mathrm{sp}}^{k}+\mathbf{r}^{k}$ where $\mathbf{x}_{\mathrm{sp}}^{k}=\sum_{i=1}^{n-k}\xi^{x}_{\mathrm{sp},i}\mathbf{u}_{\mathrm{sp},i}$ and $\mathbf{r}^{k}=\sum_{i=n-k+1}^{n}\xi^{x}_{\mathrm{sp},i}\mathbf{u}_{\mathrm{sp},i}$ with $\mathbf{u}_{\mathrm{sp},i}$ the $i$th spatial eigenvector, and (ii) replace $\mathbf{x}$ by the unconfounded component $\mathbf{r}^{k}$ in model (3). Proposition 1 shows that the bias of the corresponding covariate effect estimate is $\langle\mathbf{r}^{k},\mathbf{z}\rangle_{\bm{\mathbf{\Sigma}}^{-1}}/\langle\mathbf{r}^{k},\mathbf{r}^{k}\rangle_{\bm{\mathbf{\Sigma}}^{-1}}$ which should therefore, in theory, be 00 due to the assumption of unconfoundedness. However, as the component $\mathbf{r}^{k}$ can also be modelled as part of the highest frequencies of the spatial effect in (3), the estimate of $\beta$ obtained in this way is highly sensitive to inaccuracies at these frequencies of the estimated spatial effect. Such inaccuracies easily arise in practice as the estimation is based on overall spatial smoothing, driven by the fitted values rather than the behaviour at individual spatial frequencies. In order to avoid such sensitivity, we adapt the model matrix of the regression in (ii) by removing the columns corresponding to the frequencies above $k$. This more explicitly reflects the assumption of unconfoundedness at these frequencies but allows the spatial effect to continue to adjust for confounders at frequencies below the cap. We call the estimate of $\beta$ obtained from (i) and the adapted (ii) the capped spatial+ estimate. By construction, if the assumption of unconfoundedness above the cap is true, the bias for this estimate is 00. In practice, the true value of the cap $k$ is unlikely to be known, and it may therefore be mis-specified at a value $k^{\prime}\neq k$. If $k^{\prime}<k$, $\mathbf{r}^{k^{\prime}}=\sum_{i=n-k^{\prime}+1}^{n}\xi^{x}_{\mathrm{sp},i}\mathbf{u}_{\mathrm{sp},i}$ is simply a smaller component of $\mathbf{r}^{k}$ which is still orthogonal to the spatial part of the adjusted spatial model. Therefore, the bias should still be eliminated. However, if $k^{\prime}>k$, then $\mathbf{r}^{k^{\prime}}$ contains a potentially confounded component of $\mathbf{x}$, namely, the $k^{\prime}-k$ highest frequencies of $\mathbf{x}_{\mathrm{sp}}^{k}$, and by Corollary 1 the bias becomes $\sum_{i=n-k^{\prime}+1}^{n-k}\xi^{x}_{\mathrm{sp},i}\xi^{z}_{\mathrm{sp},i}w_{i}/\sum_{i=n-k^{\prime}+1}^{n}(\xi^{x}_{\mathrm{sp},i})^{2}w_{i}$. For comparison, the bias in the original spatial model (3) in this scenario is $\sum_{i=1}^{n-k}\xi^{x}_{\mathrm{sp},i}\xi^{z}_{\mathrm{sp},i}w_{i}/\sum_{i=1}^{n}(\xi^{x}_{\mathrm{sp},i})^{2}w_{i}$. Thus, the numerator for the spatial model includes additional confounded frequencies contributing to the bias, but at the same time the denominator is also larger. Therefore, in this case, the capped spatial+ estimate may not reduce and could even increase the size of the bias of the spatial model.
For a given data set in the Case 2 scenario of Section 5.2, under the assumption of unconfoundedness at the highest frequencies, the bias can therefore be assessed using capped spatial+ with a range of caps $k^{\prime}$. Starting at small values of $k^{\prime}$, the corresponding estimate of $\beta$ should be unbiased and therefore stay stable as we increase $k^{\prime}$, but when $k^{\prime}$ reaches the true cap $k$, the estimate changes due to the introduction of confounding bias. Thus, if we observe stable capped spatial+ estimates for caps below a certain value, these estimates can be expected to be unbiased.
Finally, we note that the assumption of unconfoundedness does not have to be limited to the highest frequencies of the spatial effect. If we believe the unconfounded behaviour of $\mathbf{x}$ is in a different frequency range, we can easily adapt the capped spatial+ method by replacing the component $\mathbf{r}^{k}$ by the part of $\mathbf{x}$ that is believed to be unconfounded.
5.4 Simulations for capped spatial+
In this section, we use simulations to illustrate the ideas behind the capped spatial+ method proposed in Section 5.3. For a fixed value of the cap $k$ we simulate data in which the covariate $\mathbf{x}$ is fully spatial but the $k$ highest frequencies are unconfounded. We then consider the bias in the spatial model (3) as well as the capped spatial+ estimates for different values of the $\mathrm{cap}$.
The data is generated such that:
$\mathbf{y}=\beta\mathbf{x}+\mathbf{z}+\bm{\epsilon}^{y},\quad\mathbf{x}=\mathbf{z}^{x}_{\mathrm{sp,low}}+\mathbf{z}^{x}_{\mathrm{sp,medium}}+a\ \mathbf{z}^{x}_{\mathrm{sp,high}},\quad\mathbf{z}=\mathbf{z}^{z}_{\mathrm{sp,low}}+\mathbf{z}^{x}_{\mathrm{sp,medium}}$
where $\bm{\epsilon}^{y}\sim N(\mathbf{0},\sigma^{2}\mathbf{I})$ and $a$ measures the weight of the unconfounded high frequencies in $\mathbf{x}$. In order to generate $\mathbf{z}^{x}_{\mathrm{sp,low}}$, $\mathbf{z}^{x}_{\mathrm{sp,high}}$ and $\mathbf{z}^{x}_{\mathrm{sp,medium}}$, we fit a thin plate spline to a realization of a Gaussian process with $\kappa=0.5/3$ (see beginning of the section), where the penalty and design matrix have now been reparameterised such that they only consider the $100$ lowest, the $k$ highest and all medium frequencies, respectively; $\mathbf{z}^{z}_{\mathrm{sp,low}}$ is generated similarly to Section 4 but it is fitted to a new realization of a Gaussian process with $\kappa=0.5/3$ using only the $100$ lowest basis functions. We consider $k\in\{5,10,20\}$ and $a\in\{1,2\}$ in the data generating process; $\beta=1$, $\sigma=0.1$, $n=1000$, $N=50$. The reparametrisation is explained in Web Appendix 2.
In the data analysis model, the capped spatial+ model uses a reparameterised $\tilde{\mathbf{B}}_{\mathrm{sp}}$ and $\tilde{\bm{\beta}}_{\mathrm{sp}}$ in the design and penalty matrix for the spatial effects in the first and second stage equation of spatial+ with $\mathrm{cap}\in\{5,10,15,20,25,30,40,50\}$, so that we consider cases where the used cap is above, below or equal to the true cap $k$. Unlike the standard spatial+, this allows us to control the frequencies included in the spatial effects for $\mathbf{y}$ and $\mathbf{x}$, such that ideally $\mathbf{r}^{x}$ should not have any confounded frequencies. Crucially here, on the first stage equation for $\mathbf{x}$ we use no smoothing in order to guarantee the residuals $\mathbf{r}^{x}$ are orthogonal to the subspace spanned by $\tilde{\mathbf{B}}_{\mathrm{sp}}\tilde{\bm{\beta}}_{\mathrm{sp}}$. We use $n$ basis functions in the spatial model.
In Figure 3, the bias decreases for increasing $k$ and $a$, as larger $k$ implies less confounded medium frequencies and more unconfounded high frequencies, and larger $a$ implies more weight being put on unconfounded high frequencies in $\mathbf{x}$ and thus less bias. For $\mathrm{cap}\leq k$, we are able to completely remove confounding bias, even if only using a smaller fraction of the unconfounded high frequency component in $\mathbf{x}$ to identify the effect. When the $k<\mathrm{cap}$ the residuals in spatial+ contain medium frequency confounding and, quite noticeably, bias arises. Thus, we observe a constant bias of zero for caps below $k$ and a considerable increase for caps above $k$. The MSE for the predicted values reveals that in general capping increases MSE, although quite negligibly for $k\leq\mathrm{cap}$, when compared to the drastic increase observed for $k>\mathrm{cap}$ (see Web Appendix 3.1).
Figure 3: Bias of $\widehat{\beta}$ in spatial model and capped spatial+ models with different combinations of $k\in\{5,10,20\}$ and $a\in\{1,2\}$. The true value is $\widehat{\beta}=1$. All other parameters remain constant.
6 Application to air temperature in Germany
As a motivating study, we analyse the relationship between the monthly mean temperature at two metres above ground, treated here as the response, and the monthly mean precipitation, the covariate, in Germany for the year 2010 at 336 locations. The data are open access and provided by the German Meteorological Service (DWD). For convenience, precipitation is measured in units of 10 millimetres and temperature is measured in degrees Celsius. In the context of spatial confounding, climate is a compelling example, as both covariates and response can exhibit relatively larger or narrower spatial correlation ranges, also depending on the time of year.
Independent regression models are fitted to each month, where rainfall and temperature are first standardised and the resulting regression coefficient for rainfall is converted back to the original scale. We consider regression models: (1) capped spatial+ (see Section 5.3), (2) spatial, (3) non-spatial. By comparing the results of (1) and (2), we show how capped spatial+ can help identify problematic covariate effect estimates in the spatial model and suggest appropriate adjustments.
All spatial effects are modelled with thin plate spline based on mgcv. The results are shown in Figure 4. As the cap used in our spatial+ increases, the fit of the model deteriorates, so we only show the estimates with $cap=5,6,\ldots,15$ (caps below 5 may not be reliable as there is likely little spatial information left in the residuals), which are also statistically significant at the 95% confidence level. From September to January and May, the caps (below 15) are generally insignificant, as indicated by the lack of pink dots in the plot. In the other months, the capped spatial+ can be used to assess whether we have confidence in the spatial estimate, and the dots show how the estimates should likely be adjusted. November is perhaps the most puzzling, where the coefficients appear to be either positive or negative depending on the frequencies included in the capped spatial+. On closer inspection, this can be explained by a more complex spatial structure of precipitation in this month, with many localised patterns, so that a linear effect of precipitation may not be sufficient in general. With the exception of November and December, all estimates plotted in Figure 4 are negative, indicating a generally negative association between the two variables.
Figure 4: Coefficient estimates of $\hat{\beta}$. The solid lines correspond to the spatial model (teal), the non-spatial model (grey) and a coefficient estimate of zero (black). The pink dots show the statistically significant at 5% estimates of the capped spatial+ for the first 15 caps. The triangles show the first two statistically significant results from capped spatial+ using the sliding window.
In order to understand the observed behaviour in more detail, we focus on three months: January, April and August. The results are shown in Table 1 for caps up to 9, which also shows the AIC and RMSE of the different models. We start with the spring and summer months. In April, the capped spatial+ estimates are lower in absolute value than those of the spatial model. This may indicate that the spatial model estimate is biased, and under the assumption that the highest frequencies are unconfounded, the estimate should be adjusted down (in absolute terms). In August, the estimates for all models appear to be fairly consistent, with little variability in the capped spatial+ estimates. This suggests that unobserved spatial confounders are not affecting the estimation and the spatial model estimate does not need to be adjusted. In January, we know from Figure  that none of the capped spatial+ estimates up to a cap of 15 are statistically significant. This suggests that the highest frequencies (up to 15) in the spatial effect are not helping us to identify the effect of rainfall and may simply not be present in this particular covariate. As mentioned in Section 5.2, however, when a covariate is dominated by low frequencies, the estimate in the spatial model may well be biased. In this case, we can adapt the capped spatial+ method to identify the highest frequencies at which $\mathbf{x}$ has a significant effect. Assuming that these frequencies are unconfounded, we can use the corresponding capped spatial+ estimate to assess the bias. We use a sliding window technique that systematically removes 15 consecutive frequencies, starting progressively deeper into the low frequency spectrum, where $start$ is the starting frequency. Figure 4 shows the estimates for the first two sliding windows, with significant estimates at the 5% level represented by triangles. We can see that in January the spatial model may overestimate the effect (in absolute terms). We followed the same approach for all the remaining autumn and winter months and for May. The results suggest that the extra peak building up to October may be an artefact, with values likely closer to flat around $-0.125$. Apart from the peak in April – perhaps a case of “Der April macht, was er will” (“April does as it pleases”) – the coefficients remain relatively stable. In all the months analysed using the sliding window technique, the first significant coefficient appears the latest at $start=8$, except in September, where it appears at $start=38$.
Table 1: Coefficient estimates for the regression models fitted to average monthly air temperature.
month model/cap $\hat{\beta}$ p-value AIC RMSE
1 non-spatial -0.0922849 0.0524941 1306.6419 0.9897
1 spatial -0.3885536 0.0000000 650.9834 0.1441
1 cap =5 -0.1847219 0.1679208 870.5588 0.2706
1 cap =6 0.0654211 0.5980586 893.6362 0.2911
1 cap =7 0.1690087 0.1617385 903.5738 0.3013
1 cap =8 0.1773077 0.1480378 915.2859 0.3124
1 cap =9 0.0240164 0.7846951 917.0572 0.3159
4 non-spatial -0.5811757 0.0000000 1196.6612 0.7748
4 spatial -0.7732586 0.0000000 951.3315 0.3053
4 cap = 5 -0.6189845 0.0000023 1095.9985 0.4349
4 cap = 6 -0.5906941 0.0000051 1096.8946 0.4380
4 cap = 7 -0.6318752 0.0000005 1098.1054 0.4404
4 cap = 8 -0.6166930 0.0000002 1096.3102 0.4387
4 cap = 9 -0.3856196 0.0013258 1148.0541 0.5282
8 non-spatial -0.1244348 0.0000000 1231.8161 0.8173
8 spatial -0.1292622 0.0000000 1012.2477 0.3297
8 cap =5 -0.1382547 0.0000314 1157.0444 0.5120
8 cap =6 -0.1262546 0.0000555 1156.6078 0.5134
8 cap =7 -0.1248357 0.0000148 1155.2212 0.5126
8 cap =8 -0.1270226 0.0000038 1153.4515 0.5106
8 cap =9 -0.1530327 0.0000000 1154.1968 0.5184
7 Discussion
As illustrated in this paper, the analytical expressions derived in Proposition 1 and Corollary 1 provide an intuitive framework for explaining and, indeed, demystifying the bias arising from spatial confounding. In particular, we show that bias in the spatial model is caused by spatial smoothing and can become arbitrarily large. In practical applications, the difference between the estimates in the spatial and non-spatial models is commonly seen as an indicator of spatial confounding. However, our analysis and simulations show that this difference is not directly related to the size or occurrence of bias in the spatial model. Instead we propose using the methods set out in Section 5 to assess and adjust for the bias.
As most other work on spatial confounding, the focus in this paper has been on quantifying the bias in spatial regressions. A derivation and short discussion about the variance of the covariate effect estimate is included in Web Appendix 1.6. However, to put the bias into perspective and thoroughly conduct statistical inference, more investigations on associated uncertainty will be an important direction for future research.
Data availability statement
The data is publicly available on the website of the German weather service (DWD).
Supporting Information
Web Appendices referenced in Section 26, are included below.
See pages - of appendix.pdf

### v3

@@
 The problem is complex and has been the topic of extensive research with sometimes puzzling and seemingly contradictory results.
-Here, we develop a broad theoretical framework that brings mathematical clarity to the mechanisms of spatial confounding, providing explicit analytical expressions for the resulting bias.
+Here, we develop a broad theoretical framework that brings mathematical clarity to the mechanisms of spatial confounding, deriving an explicit analytical expression for the resulting bias.
 We see that the problem is directly linked to spatial smoothing and identify exactly how the size and occurrence of bias relate to the features of the spatial model as well as the underlying confounding scenario.
@@
 We illustrate our approach with an application to air temperature in Germany.
-Keywords: Confounding bias, spatial regression, spatial random effects, smoothing, spatial+.
+Keywords: Confounding bias; spatial regression; spatial random effects; smoothing; spatial+.
 1 Introduction
@@
 $y(s)=\beta x(s)+\gamma(s)+\epsilon(s)$ (1)
-where $y(s)$ denotes the response variable of interest, $x(s)$ a covariate with regression coefficient $\beta$, $\gamma(s)$ a spatially correlated random effect and $\epsilon(s)$ independent identically distributed error terms, all evaluated at locations $s$ within some spatial domain.
-We consider the case of only one covariate but the model could be expanded to include more.
+where $y(s)$ denotes the response variable of interest, $x(s)$ a covariate with regression coefficient $\beta$, $\gamma(s)$ a spatially correlated random effect and $\epsilon(s)$ independent and identically distributed error terms, all evaluated at locations $s$ within some spatial domain.
 The spatial effect $\gamma(s)$ is included to approximate unmeasured spatial variation in the response which typically arises due to unobserved spatial variables.
@@
 This implies independence between the covariate and the spatial effect, which is unlikely to hold for observational data where covariates are realisations of potentially spatially dependent random variables.
-If indeed this assumption is violated, it leads to a phenomenon known as spatial confounding where the estimate of the covariate effect $\beta$ becomes biased even when the model is correctly specified.
-Thus, understanding spatial confounding is fundamental to reliable covariate effect estimation in spatial statistics, and the problem has received considerable attention in the literature since the seminal work of Clayton et al., (1993) and, later, Hodges and Reich, (2010).
+If indeed this assumption is violated, it leads to a phenomenon known as spatial confounding, where the estimate of the covariate effect $\beta$ becomes biased even when the model is correctly specified.
+Thus, understanding spatial confounding is fundamental to reliable estimation of covariate effects in spatial statistics, and the problem has received considerable attention in the literature since the seminal work of Clayton et al., (1993) and, later, Hodges and Reich, (2010).
 Recently, there has been a surge in interest, not least because the established understanding and methods most commonly used for dealing with the problem have been shown to be problematic.
 Broadly, research into spatial confounding can be structured along the following lines: (i) Investigations and suggestions for potential solutions under specific instances of the model (1) such as a choice of discrete or continuous spatial dependence and/or a particular representation of the spatial effect (Dupont et al.,, 2022; Guan et al.,, 2023; Marques et al.,, 2022; Urdangarin Iztueta et al.,, 2022).
-(ii) Identifying factors that influence the size or occurrence of bias, in particular, the importance of the spatial scales of both measured and unmeasured variables (Paciorek,, 2010; Page et al.,, 2017; Keller and Szpiro,, 2020; Nobre et al.,, 2021).
+(ii) Identifying factors that influence the size or occurrence of bias, in particular, the importance of the spatial scales of both measured and unmeasured variables (Paciorek,, 2010; Page et al.,, 2017; Keller and Szpiro,, 2020; Nobre et al.,, 2021; Bobb et al.,, 2022).
 (iii) Research based on a (sometimes implicit) preference for the estimate of $\beta$ in a non-spatial analysis (Reich et al.,, 2006; Hanks et al.,, 2015; Hughes and Haran,, 2013; Hefley et al.,, 2017; Adin et al.,, 2023; Briz-Redón,, 2023).
-This approach, usually involving a method known as restricted spatial regression, dominated the literature as well as practical implementations for many years, but has since been shown to be problematic for several reasons (Khan and Calder,, 2022; Zimmerman and Ver Hoef,, 2022).
+This approach, usually involving a method known as restricted spatial regression, dominated the literature as well as practical implementations for many years, but has since been shown to be problematic for several reasons (Khan and Calder,, 2022; Zimmerman and Ver Hoef,, 2022; Bradley,, 2024).
 (iv) Research based on causal inference approaches (Thaden and Kneib,, 2018; Papadogeorgou et al.,, 2019; Schnell and Papadogeorgou,, 2020), an overview of which is in Reich et al., (2021).
@@
 This is in part because the complexity of spatial models means that conclusions are often based on simulation studies, potentially sensitive to the specific simulation setup.
-Moreover, the research direction (iii) above has lead to a focus on the difference between the estimates in the spatial and non-spatial analyses for identifying confounding bias.
-However, as our results confirm, this is not directly related to the occurrence or otherwise of bias.
-There have also been attempts to prove that spatial confounding bias in models of the form (1) is always negligible so that the issue does not exist in practice (Khan and Berrett,, 2023), but we will see that bias from spatial confounding can in fact become arbitrarily large.
+Moreover, the research direction (iii) above has led to a focus on the difference between the estimates in the spatial and non-spatial analyses for identifying confounding bias.
+However, as our results confirm, this difference does not in itself determine whether either estimate is biased.
 We develop a theoretical framework that brings mathematical clarity to the mechanisms of spatial confounding and the methods used for dealing with the resulting bias.
-We see that it is a problem for spatial models in general, directly linked to spatial smoothing (in line with work on bias in linear mixed models, Schnell and Bose,, 2019).
+We show that it is a problem for spatial models in general, directly linked to spatial smoothing (in line with work on bias in linear mixed models, Schnell and Bose,, 2019).
 Central to our findings is the analytical expression for the bias derived in Proposition 1 which shows that, in the metric defined by the precision matrix of the chosen spatial analysis model, the bias is the correlation between the covariate and the confounder times the size of the confounder relative to the size of the covariate.
-Not only is this an intuitive result, but by using an eigendecomposition of the precision matrix, we can see how the bias depends on the underlying confounding scenario and the covariance structure and parameters of the analysis model.
-We also see explicitly the role of spatial frequencies, defined directly in terms of the model.
-While the expression for the bias is relatively simple, the dependence on individual components and parameters can be complex, explaining the subtle and at times counter-intuitive behaviours.
+Not only is this an intuitive result, but by using an eigendecomposition of the precision matrix, we derive a nuanced and detailed understanding of how the bias depends on the underlying confounding scenario and the covariance structure and parameters of the analysis model.
+This includes a deepened understanding of the role of spatial frequencies, defined directly in terms of the analysis model and the resulting eigendecomposition.
+While the expression for the bias is itself relatively simple to derive and aligns with earlier results on regularized regression, the dependence on individual components and parameters in the spatial context is complex, explaining the subtle and at times counter-intuitive behaviours.
 We illustrate our findings in a simulation study of several different scenarios.
+A number of recent works (Dupont et al.,, 2022; Gilbert et al.,, 2024; Bolin and Wallin,, 2024) have investigated the asymptotic behaviour of the covariate effect estimate in models of the form (1) under infill asymptotics, in particular, Gilbert et al., (2024) and Dupont et al., (2022) show that if the covariate has non-spatial information, the bias tends to zero as the sample size goes to infinity.
+Of course, in practice, this limit is never reached.
+Our focus in this paper is on the behaviour of the covariate effect estimate in (1) for a finite sample size.
+We show that, in this case, there are scenarios in which spatial confounding bias is negligible (in line with Khan and Berrett, (2023)), however, there are also scenarios in which the bias can become arbitrarily large.
 Based on our results, we develop a general framework for assessing and adjusting spatial confounding bias in practice, applicable under any specification of the spatial effects in (1).
@@
 Specifically, we propose a procedure in which multiple capped versions of spatial+ are applied.
+The method can easily be adapted to the scenario in which other frequencies than the highest ones are assumed to be unconfounded.
 We illustrate our approach with an application to monthly mean air temperature across Germany during the year 2010 with rainfall as covariate.
@@
 We show how our methods can be applied to assess whether the effect estimate in the spatial model is biased and, if so, determine the estimated bias.
+In the following, we first establish some preliminary results on spatial regression models in Section 2 that will allow us to derive compact and intuitive expressions for bias arising due to spatial confounding in Section 3, supplemented by detailed simulations in Section 4.
+In Section 5, we propose a method for assessing bias in practice, supplemented by an illustrative simulation study.
+Finally, in Section 6, we illustrate our approach with an application.
 2 Definitions and preliminary results
@@
 where $\mathbf{y}=(y(s_{1}),\ldots,y(s_{n}))^{T}$, $\mathbf{x}=(x(s_{1}),\ldots,x(s_{n}))^{T}$ and $\bm{\epsilon}=(\epsilon(s_{1}),\ldots,\epsilon(s_{n}))^{T}$ are the response, covariate and error term at locations $s_{1},\ldots,s_{n}$ and $\mathbf{z}=(z(s_{1}),\ldots,z(s_{n}))^{T}$ the true confounding variable at these locations.
-Since the confounder is assumed to be unknown, in the spatial analysis model, chosen by the researcher analysing the data, we assume that $\mathbf{z}(s)$ is approximated by a spatial effect $\gamma(s)$ as in (1).
+Since the confounder is assumed to be unknown, in the spatial analysis model chosen by the researcher analysing the data, we assume that $\mathbf{z}(s)$ is approximated by a spatial effect $\gamma(s)$ as in (1).
 This implies an analysis model of the form
@@
 Owing to the generality of our setup, all the commonly used spatial models fit into this generic framework.
-For example, for (i) Gaussian random fields based on coordinate information, $\mathbf{B}_{\mathrm{sp}}$ is an incidence matrix linking observations to observation locations and $\mathbf{S}$ is the inverse covariance matrix implied by the covariance function (ii) (intrinsic) conditional auto-regressive and Gaussian Markov random field specifications for regional data, $\mathbf{B}_{\mathrm{sp}}$ is an indicator matrix for the regions and $\mathbf{S}$ encodes spatial proximity with an adjacency matrix or graph Laplacian, and (ii) tensor product P-splines and thin plate splines, $\mathbf{B}_{\mathrm{sp}}$ contains basis function evaluations and $\mathbf{S}$ the corresponding smoothness penalty.
+For example, for (i) Gaussian random fields based on coordinate information, $\mathbf{B}_{\mathrm{sp}}$ is an incidence matrix linking observations to observation locations and $\mathbf{S}$ is the inverse covariance matrix implied by the covariance function (ii) (intrinsic) conditional auto-regressive and Gaussian Markov random field specifications for regional data, $\mathbf{B}_{\mathrm{sp}}$ is an indicator matrix for the regions and $\mathbf{S}$ encodes spatial proximity with an adjacency matrix or graph Laplacian, and (iii) tensor product P-splines and thin plate splines, $\mathbf{B}_{\mathrm{sp}}$ contains basis function evaluations and $\mathbf{S}$ the corresponding smoothness penalty.
 We can also take a Bayesian perspective if we have a multivariate Gaussian prior on $\bm{\beta}_{\mathrm{sp}}$ with precision $\lambda\mathbf{S}$, see Fahrmeir and Kneib, (2011) for details on both the various options for spatial smoothing and the Bayesian perspective on spatial regression.
@@
 Central to our further analyses are the properties of the precision matrix $\bm{\mathbf{\Sigma}}^{-1}$ of the spatial analysis model (3), derived from the covariance structure $\bm{\mathbf{\Sigma}}=\sigma^{2}\mathbf{I}+\lambda^{-1}\mathbf{B}_{\mathrm{sp}}\mathbf{S}^{-}\mathbf{B}_{\mathrm{sp}}^{T}$.
-Details of our derivations are included in Web Appendix 1.1.
-Lemma 1.
+Details of our derivations are included in Appendix 1.1.
+The following lemma shows how the eigenvectors of $\bm{\mathbf{\Sigma}}^{-1}$ are linked to the penalisation (i.e.
+the spatial smoothing) in the chosen analysis model.
+Lemma 1
 Let $\alpha_{1}\leq\cdots\leq\alpha_{p}$ be the eigenvalues of the penalty matrix $\mathbf{S}$ and $\lambda>0$ the smoothing parameter.
 Then the eigenvalues of the precision matrix $\bm{\mathbf{\Sigma}}^{-1}$ are given by $\{\sigma^{-2},\sigma^{-2}w_{1},\ldots,\sigma^{-2}w_{p}\}$ where $w_{i}=\lambda\alpha_{i}/(\sigma^{-2}+\lambda\alpha_{i})\textrm{ for }i=1,\ldots,p.$
-Lemma 1 shows that there exists an $n\times n$-matrix $\mathbf{U}$ such that $\mathbf{U}^{T}\bm{\mathbf{\Sigma}}^{-1}\mathbf{U}=\sigma^{-2}\textrm{diag}(1,\ldots,1,w_{1},\allowbreak\ldots,w_{p})$ where $w_{i}=\lambda\alpha_{i}/(\sigma^{-2}+\lambda\alpha_{i})$ for $i=1,\dots,p$.
-The columns of $\mathbf{U}$ are an orthonormal basis of eigenvectors.
-The first $n-p$ eigenvectors $\mathbf{U}_{\mathrm{ns}}$ span the non-spatial part of $\mathbb{R}^{n}$ which is unaffected by smoothing.
-The remaining $p$ eigenvectors $\mathbf{U}_{\mathrm{sp}}$ represent different types of spatial behaviour, where the $i$th spatial eigenvector has eigenvalue $\sigma^{-2}w_{i}$ with $w_{i}$ a weight between 00 and $1$, directly linked to the smoothing penalty $\alpha_{i}$.
-Since $w_{1}\leq\cdots\leq w_{p}$, it is the spatial vectors with higher weights that correspond to behaviours that are penalised more, i.e.
-that are considered “less spatial” according to the analysing model.
-Thus, we will refer to behaviour given by a spatial eigenvector with weight $w_{i}$ close to 1 as being “high frequency” and eigenvectors with low weights $w_{i}$ as “low frequency”.
-In this way, what is considered high/low frequency spatial behaviour is always well-defined and directly linked to the spatial structure chosen for the spatial analysis model.
-Since the precision matrix $\bm{\mathbf{\Sigma}}^{-1}$ is symmetric positive definite it defines an inner product on $\mathbb{R}^{n}$.
+Using this result, we obtain a natural decomposition of the sample space $\mathbb{R}^{n}$. Specifically, Lemma 1 shows that there exists an $n\times n$-matrix $\mathbf{U}$ such that
+$\mathbf{U}^{T}\bm{\mathbf{\Sigma}}^{-1}\mathbf{U}=\sigma^{-2}\textrm{diag}(1,\ldots,1,w_{1},\allowbreak\ldots,w_{p}),$
+where $w_{i}=\lambda\alpha_{i}/(\sigma^{-2}+\lambda\alpha_{i})$ for $i=1,\dots,p$, and the columns of $\mathbf{U}$ form an orthonormal basis of eigenvectors in $\mathbb{R}^{n}$.
+As elements of the sample space, each of the eigenvectors represent different behaviours over the spatial domain.
+The first $n-p$ eigenvectors $\mathbf{U}_{\mathrm{ns}}$ have eigenvalue $\sigma^{-2}$ and are unaffected by smoothing.
+Therefore, the subspace spanned by these can be considered as the “non-spatial” part of the sample space.
+The remaining $p$ eigenvectors $\mathbf{U}_{\mathrm{sp}}$ represent the spatially smoothed behaviours, where the $i$th eigenvector has eigenvalue $\sigma^{-2}w_{i}$ with $w_{i}$ a weight between 00 and $1$, directly linked to the $i$th smoothing penalty $\alpha_{i}$.
+Although, the values of the weights depend on the parameters $\sigma^{-2}$ and $\lambda$, which are usually estimated, the eigenvectors do not.
+Moreover, irrespective of the parameter estimates, the ordering of the weights is the same as the $\alpha_{i}$’s, i.e.
+$0\leq w_{1}\leq\cdots\leq w_{p}\leq 1$
+Thus, spatial eigenvectors with weights close to $1$ correspond to behaviours that are penalised more, i.e.
+that are considered “less spatial” according to the analysis model.
+We will therefore refer to behaviours given by eigenvectors with high weights $w_{i}$ as “high frequency” and eigenvectors with low weights $w_{i}$ as “low frequency”.
+For further details, see Appendix 1.2.
+Although spatial frequencies are generally known to play an important role for spatial confounding, there is no single definition in the literature for what is meant by spatial frequencies (see Appendix 1.2).
+Describing them directly through the eigenvectors of the spatial precision matrix $\bm{\mathbf{\Sigma}}^{-1}$ has several advantages:
+  • $\bullet$
+    For any choice of analysis model (3), these eigenvectors are well-defined and can be easily computed, even when the dimension $p$ of the spatial effect is smaller than the sample size $n$;
+  • $\bullet$
+    The eigenvectors form an orthonormal basis for the $n$-dimensional sample space;
+  • $\bullet$
+    The eigenvectors are directly and quantifiably linked to the penalisation (i.e. spatial smoothing) in the chosen analysis model (as set out in Lemma 1).
+Thus, what is considered high/low frequency spatial behaviour is always well-defined and determined by the spatial structure chosen for the spatial analysis model.
+Finally, since the precision matrix $\bm{\mathbf{\Sigma}}^{-1}$ is symmetric positive definite it defines an inner product on $\mathbb{R}^{n}$.
 For the rest of this paper, if $\mathbf{M}$ is a symmetric positive definite $n\times n$ matrix, we will use the notation $\langle\cdot,\cdot\rangle_{\mathbf{M}}$ to denote the inner product defined by $\mathbf{M}$ and $\|\cdot\|_{\mathbf{M}}$ the corresponding norm.
@@
 3 Characterising the impact of spatial confounding
-In this section we present the central results of the paper, namely, Proposition 1 and Corollary 1, which provide explicit expressions for the bias of the covariate effect estimate in the analysis model (3) arising from spatial confounding.
+In this section, Proposition 1 and Corollary 1 provide explicit expressions for the bias of the covariate effect estimate in the analysis model (3) arising from spatial confounding.
 We assume throughout that the true data generating process is that of model (2) with true covariate effect $\beta$ and true spatial confounder $\mathbf{z}$, where the latter is assumed to be generated from the spatial structure of model (3).
@@
 3.1 Bias in the spatial model
-Proposition 1.
+Proposition 1
 The bias of the estimated covariate effect $\widehat{\beta}$ in model (3) is given by
@@
 the sign of $\langle\mathbf{x},\mathbf{z}\rangle_{\bm{\mathbf{\Sigma}}^{-1}}$.
-The proof of the proposition is in the appendix.
+The proof of the proposition is in Appendix 1.3.
 Combining Proposition 1 with the results of Section 2 we get a better understanding of how the bias behaves.
@@
 Since $\mathbf{z}$ is spatial, $\bm{\xi}^{z}_{\mathrm{ns}}=\mathbf{0}$.
-It is now straightforward to prove the following result (see the Web Appendix for the proof).
-Corollary 1.
+It is now straightforward to prove the following result (see Appendix 1.4.
+for the proof).
+Corollary 1
 Let $\mathbf{U}$ be the orthonormal eigenbasis which diagonalises $\bm{\mathbf{\Sigma}}^{-1}$, and $\bm{\xi}^{x}=(\bm{\xi}^{xT}_{\mathrm{ns}},\bm{\xi}^{xT}_{\mathrm{sp}})^{T}$ and $\bm{\xi}^{z}=(\mathbf{0}^{T},\bm{\xi}^{zT}_{\mathrm{sp}})^{T}$ the coordinates of $\mathbf{x}$ and $\mathbf{z}$ in this basis.
@@
 The expression for the bias is not symmetric in $\mathbf{x}$ and $\mathbf{z}$.
-In particular, it is not affected by components in $\mathbf{z}$, i.e.
-with $\xi_{\mathrm{sp},i}^{z}\neq 0$, for which the corresponding component in $\mathbf{x}$ is 0.
+In particular, it is not affected by components in $\mathbf{z}$ for which the corresponding component in $\mathbf{x}$ is 0 (i.e.
+when $\xi_{\mathrm{sp},i}^{z}\neq 0$ but $\xi_{\mathrm{sp},i}^{x}=0$).
 Thus, any unmeasured spatial variation in the response variable that is independent of the covariate can be ignored for the purposes of the bias.
@@
 Due to the differences in the weights $w_{i}$, different spatial frequencies affect the bias differently.
-Looking at the numerator in Corollary 1, the larger the weight, the larger the contribution to the bias.
-So confounding at high frequencies (where weights are close to $1$) induce larger bias, while confounding at low frequencies may give negligible contribution to the bias.
+Looking at the numerator in Corollary 1, any confounded spatial component of $\mathbf{x}$ (i.e.
+those with $\xi_{\mathrm{sp},i}^{x}\neq 0$ and $\xi_{\mathrm{sp},i}^{z}\neq 0$ ) contribute to the bias, and the larger the weight, the larger the contribution.
+So confounding at high frequencies (where weights are close to $1$) induce larger bias, while confounding at low frequencies may give a negligible contribution.
 Confounding at unpenalised spatial frequencies does not lead to bias (as they have weights 00), and non-spatial components in $\mathbf{x}$ do not contribute to the numerator either.
@@
 Nevertheless, we consider how the bias in model (4) compares to the spatial model (3) as this has been a focus of spatial confounding research in the past.
-The following corollary is proved in the appendix.
-Corollary 2.
+The following corollary is proved in Appendix 1.8.
+Corollary 2
 Let $\mathbf{U}$ be the orthonormal eigenbasis which diagonalises $\bm{\mathbf{\Sigma}}^{-1}$, and $\bm{\xi}^{x}=(\bm{\xi}^{xT}_{\mathrm{ns}},\bm{\xi}^{xT}_{\mathrm{sp}})^{T}$ and $\bm{\xi}^{z}=(\mathbf{0}^{T},\bm{\xi}^{zT}_{\mathrm{sp}})^{T}$ the coordinates of $\mathbf{x}$ and $\mathbf{z}$ in this basis.
@@
 However, as the weights affect both the numerator and denominator of the bias, in general, the relationship between the two models is not straightforward.
-A detailed computation of the difference and how it is affected by different types of behaviours is given in Web Appendix 1.4.
+A detailed computation of the difference and how it is affected by different types of behaviours is given in Appendix 1.4.
 Although the bias tends to be larger in the non-spatial model than in the spatial model, there are scenarios where it is the other way around.
 In the past, the estimate in model (4) was often mistakenly assumed to be unbiased and, therefore, a large difference between this and the estimate in (3) was seen as a sign of confounding bias in the spatial model (and, conversely, a small difference a sign of no confounding).
-However, Corollary 2 shows that a difference between these estimates tends to arise when the covariate is dominated by low frequencies (as these have weights most different from 1) and is not directly related to the existence or otherwise of confounding bias.
+However, Corollary 2 shows that a difference between these estimates tends to arise when the covariate is dominated by low frequencies (as these have weights most different from $1$) and the existence of such a difference does not necessarily mean that there is confounding bias in the spatial model.
+Conversely, we see that if the covariate is dominated by high frequencies (with weights close to $1$), the estimates in the two models will be similar, but this does not necessarily mean that they are unbiased.
 3.4 Dependence on the smoothing parameter
@@
 However, as $\lambda$ affects both the numerator and the denominator of the expression in Corollary 1, there can be situations where this is not the case.
-In Web Appendix 1.5 we provide a detailed analysis which shows that, although the behaviour of the bias varies, when there is sufficient non-spatial or unconfounded high frequency components in $\mathbf{x}$, the size of the bias broadly increases with $\lambda$ (in line with intuition).
-The following corollary is also proved in the Web Appendix.
-Corollary 3.
+In Appendix 1.5 we provide a detailed analysis which shows that, although the behaviour of the bias varies, when there is sufficient non-spatial or unconfounded high frequency components in $\mathbf{x}$, the size of the bias broadly increases with $\lambda$ (in line with intuition).
+The following corollary is also proved in the Appendix 1.6.
+Corollary 3
 Let $\mathbf{U}$ be the orthonormal eigenbasis which diagonalises $\bm{\mathbf{\Sigma}}^{-1}$, and $\bm{\xi}^{x}=(\bm{\xi}^{xT}_{\mathrm{ns}},\bm{\xi}^{xT}_{\mathrm{sp}})^{T}$ and $\bm{\xi}^{z}=(\mathbf{0}^{T},\bm{\xi}^{zT}_{\mathrm{sp}})^{T}$ the coordinates of $\mathbf{x}$ and $\mathbf{z}$ in this basis.
@@
 where $\beta\in\mathbb{R}$, $\bm{\epsilon}^{y}\sim N(\mathbf{0},\sigma^{2}\mathbf{I})$, $\bm{\epsilon}^{x}\sim N(\mathbf{0},\sigma^{2}_{x}\mathbf{I})$ for $\sigma,\sigma_{x}>0$.
-In this specification, $\bm{\epsilon}^{x}$ represents the non-spatial component of $\mathbf{x}$, while $\mathbf{z}$ and $\mathbf{z}^{x}$ are constructed as linear combinations of two spatial fields $\mathbf{z}_{\mathrm{sp,l}}$ and $\mathbf{z}_{\mathrm{sp,h}}$, which are low and high frequency, respectively, i.e.
-$\mathbf{z}=\xi_{\mathrm{sp,l}}^{z}\mathbf{z}_{\mathrm{sp,l}}+\xi_{\mathrm{sp,h}}^{z}\mathbf{z}_{\mathrm{sp,h}}$ and $\mathbf{z}^{x}=\xi_{\mathrm{sp,l}}^{x}\mathbf{z}_{\mathrm{sp,l}}+\xi_{\mathrm{sp,h}}^{x}\mathbf{z}_{\mathrm{sp,h}}$ for $\xi_{\mathrm{sp,l}}^{z},\xi_{\mathrm{sp,h}}^{z},\xi_{\mathrm{sp,l}}^{x},\xi_{\mathrm{sp,h}}^{x}\in\mathbb{R}$.
+In this specification, $\mathbf{z}$ and $\mathbf{z}^{x}$ are constructed as linear combinations of two spatial fields $\mathbf{z}_{\mathrm{sp,l}}$ and $\mathbf{z}_{\mathrm{sp,h}}$, which are low and high frequency, respectively, i.e.
+$\mathbf{z}=\xi_{\mathrm{sp,l}}^{z}\mathbf{z}_{\mathrm{sp,l}}+\xi_{\mathrm{sp,h}}^{z}\mathbf{z}_{\mathrm{sp,h}},\qquad\mathbf{z}^{x}=\xi_{\mathrm{sp,l}}^{x}\mathbf{z}_{\mathrm{sp,l}}+\xi_{\mathrm{sp,h}}^{x}\mathbf{z}_{\mathrm{sp,h}}$
+for $\xi_{\mathrm{sp,l}}^{z},\xi_{\mathrm{sp,h}}^{z},\xi_{\mathrm{sp,l}}^{x},\xi_{\mathrm{sp,h}}^{x}\in\mathbb{R}$, while $\bm{\epsilon}^{x}$ is spatially unstructured and represents the non-spatial component of $\mathbf{x}$.
 Unless stated otherwise, we let $\beta=0.5$, $\sigma_{x}=\sigma=1$.
@@
 The data generating process matches the analysing spatial model which is therefore correctly specified.
-(In Web Appendix 4, we repeat the simulation where the data generating process is a Gaussian process, so in addition to the spatial confounding bias, there may be further mis-specification bias.) In order to generate $\mathbf{z}_{\mathrm{sp,l}}$ and $\mathbf{z}_{\mathrm{sp,h}}$, we start by considering a mean-zero Gaussian process $\bm{\gamma}$ with exponential covariance structure following $C(h)=\exp(-h/\kappa)$ such data $h=\lVert\mathbf{s}-\mathbf{s}^{\prime}\rVert$ for $\mathbf{s},\mathbf{s}^{\prime}\in[0,1]\times[0,1]$, which is assumed for simplicity to have variance of 1.
+In Appendix 4, we repeat the simulation where the data generating process is a Gaussian process, so in addition to the spatial confounding bias, there may be further mis-specification bias.
+In order to generate $\mathbf{z}_{\mathrm{sp,l}}$ and $\mathbf{z}_{\mathrm{sp,h}}$, we start by considering a mean-zero Gaussian process $\bm{\gamma}$ with exponential covariance structure following $C(h)=\exp(-h/\kappa)$ such data $h=\lVert\mathbf{s}-\mathbf{s}^{\prime}\rVert$ for $\mathbf{s},\mathbf{s}^{\prime}\in[0,1]\times[0,1]$, which is assumed for simplicity to have variance of 1.
 We set $\kappa=0.1$ corresponding to a spatial range of approximately 0.3.
@@
 As expected, Figure 1 (top left) shows significant bias in both models under this scenario, increasing linearly in size with $\xi_{\mathrm{sp,h}}^{z}$.
-The mean squared error (MSE) of fitted values compared to the true expectation of $\mathbf{y}$ is shown in Web Appendix 3.1 and is always lower for the spatial model, indicating a better fit, which is not surprising as the spatial model is correctly specified, whereas the non-spatial model is mis-specified.
-Figure 1: Bias of $\widehat{\beta}$ and $\widehat{\beta}_{\mathrm{ns}}$ (top) in the spatial and non-spatial (ns) models under Scenario 1 (top, left), Scenario 2 (top, right), Scenario 3 (bottom, left), Scenario 4 (bottom, right).
+The mean squared error (MSE) of fitted values compared to the true expectation of $\mathbf{y}$ is shown in Appendix 3.1 and is always lower for the spatial model, indicating a better fit, which is not surprising as the spatial model is correctly specified, whereas the non-spatial model is mis-specified.
+Figure 1: Bias of $\widehat{\beta}$ and $\widehat{\beta}_{\mathrm{ns}}$ in the spatial and non-spatial (ns) models under Scenario 1 (top, left), Scenario 2 (top, right), Scenario 3 (bottom, left), Scenario 4 (bottom, right).
 The true value of $\beta$ is 0.5.
@@
 In contrast, swapping frequencies makes no difference to the non-spatial model as all frequencies have equal weight and, indeed, the behaviour for this model is the same as in Scenario 1.
-Web Appendix 3.1 confirms, as before, that the spatial model is superior in terms of fit.
+Appendix 3.1 confirms, as before, that the spatial model is superior in terms of fit.
 4.3 Scenario 3: Comparison to the non-spatial model
-In Section 3.3 we noted that difference or similarity between the estimates in the spatial and non-spatial models is not directly linked to the existence or otherwise of bias in the spatial model.
+In Section 3.3 we noted that difference or similarity between the estimates in the spatial and non-spatial models cannot in itself be used to diagnose the existence or otherwise of bias in the spatial model.
 This is illustrated in Scenarios 1 and 2 above, as in both of these scenarios the estimates in the two models differ but, in Scenario 1, the spatial model is biased whereas, in Scenario 2, it is not.
 To illustrate this point further, here we simulate two scenarios in which the estimates in the two models are similar, but in one scenario both models are biased whereas in the other they are both largely unbiased.
-Specifically, in order to generate $\mathbf{z}_{\mathrm{sp,h}}$ and $\mathbf{z}_{\mathrm{sp,l}}$, we fit a reparametrised TPRS spatial effect (see Web Appendix 2) to the Gaussian process with $\kappa=0.1$ where the corresponding 75 highest frequencies are used to generate $\mathbf{z}_{\mathrm{sp,h}}$ and the remaining frequencies to generate $\mathbf{z}_{\mathrm{sp,l}}$.
+Specifically, in order to generate $\mathbf{z}_{\mathrm{sp,h}}$ and $\mathbf{z}_{\mathrm{sp,l}}$, we fit a reparametrised TPRS spatial effect (see Appendix 2) to the Gaussian process with $\kappa=0.1$ where the corresponding 75 highest frequencies are used to generate $\mathbf{z}_{\mathrm{sp,h}}$ and the remaining frequencies to generate $\mathbf{z}_{\mathrm{sp,l}}$.
 We consider: (1) biased scenario with $(\xi_{\mathrm{sp,l}}^{x},\xi_{\mathrm{sp,h}}^{x})=(0,4)$ and $(\xi_{\mathrm{sp,l}}^{z},\xi_{\mathrm{sp,h}}^{z})=(1,1)$; (2) unbiased scenario with $(\xi_{\mathrm{sp,l}}^{x},\xi_{\mathrm{sp,h}}^{x})=(0,0.1)$ and $(\xi_{\mathrm{sp,l}}^{z},\xi_{\mathrm{sp,h}}^{z})=(1,0)$.
@@
 The results are shown in Figure 1 (bottom, left).
-Section 3.3 also showed that while the non-spatial bias tends to exceed the spatial bias, it could also be the other way around.
-In both Scenarios 1 and 2, the non-spatial bias always stays above the spatial bias.
-However, in Web Appendix 3.2, we repeat Scenario 1 with a larger proportion of unconfounded low frequency components in $\mathbf{x}$ so that the correlation between $\mathbf{x}$ and $\mathbf{z}$ is low overall, but high at high frequencies.
-As expected from our analysis, in this latter scenario, the spatial bias exceeds the non-spatial bias.
+Section 3.3 also showed that while the bias in the non-spatial model tends to exceed that of the spatial model, it could also be the other way around.
+In both Scenarios 1 and 2, the bias in the non-spatial model always stays above the spatial model bias.
+However, in Appendix 3.2, we repeat Scenario 1 with a larger proportion of unconfounded low frequency components in $\mathbf{x}$ so that the correlation between $\mathbf{x}$ and $\mathbf{z}$ is low overall, but high at high frequencies.
+As expected from our analysis, in this latter scenario, the bias in the spatial model exceeds that of the non-spatial model.
 4.4 Scenario 4: Dependence on non-spatial information
-We now study the effect on the bias of non-spatial information in $\mathbf{x}$ by increasing the parameter $\sigma_{x}$.
-As the coefficient $\xi_{\mathrm{ns}}^{x}$ appears only in the denominator of the bias in both the spatial and the non-spatial model, we expect bias in both models to reduce.
+We now study the effect on the bias of non-spatial information in the covariate $\mathbf{x}$ by increasing the parameter $\sigma_{x}$.
+As the coefficient $\xi_{\mathrm{ns}}^{x}$ appears only in the denominator of the bias in both the spatial and non-spatial models, we expect bias in both models to reduce.
 Specifically, we consider $\sigma_{x}\in\{0.3,1,3\}$.
 Other parameters remain constant at $(\xi_{\mathrm{sp,l}}^{x},\xi_{\mathrm{sp,h}}^{x})=(1,1)$, $(\xi_{\mathrm{sp,l}}^{z},\xi_{\mathrm{sp,h}}^{z})=(0,0.5)$.
-Figure 1 (bottom, right) shows that, as expected, the bias decreases for both the non-spatial and spatial model when $\sigma_{x}$ is increased.
+Figure 1 (bottom, right) shows that, as expected, the bias decreases for both models when $\sigma_{x}$ is increased, and the case $\sigma_{x}=3$ illustrates that, for a sufficiently large non-spatial component of $\mathbf{x}$, the bias in the spatial model becomes negligible.
 As before, the MSE of fitted values is always larger for the non-spatial model.
-In Section 3.3, we saw that non-spatial bias tends to exceed that of the spatial model, but if the covariate has little non-spatial information and confounding takes place at high frequencies, then it may be the other way around.
-This behaviour is observed in Figure 1 where the non-spatial bias is higher than the spatial, except in the case $\sigma_{x}=0.3$, i.e where $\mathbf{x}$ has the smallest amount of non-spatial information.
+In Section 3.3, we saw that bias in the non-spatial model tends to exceed that of the spatial model, but if the covariate has little non-spatial information and confounding takes place at high frequencies, then it may be the other way around.
+This behaviour is observed in Figure 1 where the bias in the non-spatial model is higher than that of the spatial model, except in the case $\sigma_{x}=0.3$, i.e where $\mathbf{x}$ has the smallest amount of non-spatial information.
 4.5 Scenario 5: Dependence on the smoothing parameter
@@
 At the estimated values of $\lambda$ the bias is 0.4037, 0.1804 and 0.0362, from left to right.
-For $\sigma_{x}=0.3$, the generalised cross-validation estimate is in a relatively flat area of the curve where the spatial bias exceeds the non-spatial.
-For both of the other values of $\sigma_{x}$, although the spatial bias is bounded by the non-spatial bias, the estimate $\hat{\lambda}$ is at the steep part of the curve, which means that small changes in $\hat{\lambda}$ can lead to large changes in the bias.
+For $\sigma_{x}=0.3$, the generalised cross-validation estimate is in a relatively flat area of the curve where the bias in the spatial model exceeds that of the non-spatial model.
+For both of the other values of $\sigma_{x}$, although the spatial model bias is bounded by the non-spatial model bias, the estimate $\hat{\lambda}$ is at the steep part of the curve, which means that small changes in $\hat{\lambda}$ can lead to large changes in the bias.
 Figure 2: Bias of $\widehat{\beta}$ (solid black) and $\widehat{\beta}_{\mathrm{ns}}$ (dashed black) for increasing $\lambda$ and different $\sigma_{x}$.
@@
 As the expressions in Section 3 involve the unknown confounder $\mathbf{z}$, in practical applications, we cannot simply compute the bias to assess the impact of confounding.
-As illustrated in Section 3.3 and the Scenario 3 simulations above, we also cannot use the difference between the estimates in the spatial and non-spatial models as an indicator for the existence or otherwise of bias.
-Here, we propose a general framework for bias assessment and adjustment based on our results.
+Moreover, as illustrated in Section 3.3 and the Scenario 3 simulations above, the difference between the estimates in the spatial and non-spatial models is a poor diagnostic for the existence or otherwise of bias.
+Here, instead we propose a general framework for bias assessment and adjustment based on our results.
 That is, for a given data set and choice of spatial analysis model, we propose methods for estimating the bias under different assumptions for the data generation scenario.
@@
 As outlined below, this assumption is also important in the context of bias adjustment.
-hile the work on spatial scales is well-known, the results may be somewhat counter-intuitive, and there has been a tendency to worry more about confounding at low spatial frequencies despite this being the less problematic scenario.
-This is perhaps because low frequencies are often seen as more likely to be confounded in practice, with higher frequencies expected to capture the more characteristic behaviour of the variables.
-Moreover, as seen from our comparison between the spatial vs.
-the non-spatial model, low frequency confounding can also lead to large differences in the estimates in these two models.
-Historically, such differences were interpreted as a sign of confounding with the non-spatial estimate mistakenly assumed to be unbiased.
-Our results confirm that bias in the spatial model is unlikely to be a problem when confounding is at low frequencies, as long as there is (sufficient) non-spatial or unconfounded high frequency information in the covariate.
 5.1 Case 1: $\mathbf{x}$ has non-spatial information
@@
 Spatial+ (Dupont et al.,, 2022) is a method for bias adjustment which utilises the fact that the non-spatial component $\mathbf{x}_{\mathrm{ns}}=\mathbf{U}_{\mathrm{ns}}\bm{\xi}^{x}_{\mathrm{ns}}$ of the covariate is unconfounded and can therefore be used to identify the covariate effect.
-Although originally introduced for models in which the spatial effect is a thin plate spline, the method can be formulated for any spatial analysis model as follows.
-The model (3) with no covariates is initially fitted to $\mathbf{x}$ to give the decomposition $\mathbf{x}=\mathbf{x}_{\mathrm{sp}}+\mathbf{r}^{x}$ where $\mathbf{x}_{\mathrm{sp}}$ are the fitted values (estimating the spatial part of $\mathbf{x}$) and $\mathbf{r}^{x}$ the residuals.
-The covariate effect $\beta$ is then estimated by fitting the model (3) but with $\mathbf{x}$ replaced by $\mathbf{r}^{x}$.
-From Proposition 1, the bias in spatial+ is therefore $\langle\mathbf{r}^{x},\mathbf{z}\rangle_{\bm{\mathbf{\Sigma}}^{-1}}/\langle\mathbf{r}^{x},\mathbf{r}^{x}\rangle_{\bm{\mathbf{\Sigma}}^{-1}}$.
-As the residuals $\mathbf{r}^{x}$ estimate the non-spatial component $\mathbf{x}_{\mathrm{ns}}$ which is orthogonal to $\mathbf{z}$, this shows why we can expect the bias to be eliminated.
+Although originally introduced for models in which the spatial effect is a thin plate spline, the method can be formulated for any spatial analysis model as the following two-step procedure.
+Fit the model (3) with no covariates to $\mathbf{x}$ to obtain the decomposition $\mathbf{x}=\mathbf{x}_{\mathrm{sp}}+\mathbf{r}^{x}$ where $\mathbf{x}_{\mathrm{sp}}$ are the fitted values (estimating the spatial part of $\mathbf{x}$) and $\mathbf{r}^{x}$ the residuals.
+Estimate the covariate effect $\beta$ by fitting the model (3) but with $\mathbf{x}$ replaced by $\mathbf{r}^{x}$.
+From Proposition 1, the bias in spatial+ is $\langle\mathbf{r}^{x},\mathbf{z}\rangle_{\bm{\mathbf{\Sigma}}^{-1}}/\langle\mathbf{r}^{x},\mathbf{r}^{x}\rangle_{\bm{\mathbf{\Sigma}}^{-1}}$.
+The residuals $\mathbf{r}^{x}$ estimate the component $\mathbf{x}_{\mathrm{ns}}$ which lies in the non-spatial part of the sample space.
+As this subspace is orthogonal to the spatial part containing $\mathbf{z}$, this shows why we can expect the bias to be eliminated.
 Indeed, even though $\mathbf{r}^{x}$ may contain a small proportion of potentially confounded spatial components, its non-spatial part is likely to dominate so that bias is negligible (as seen in the Scenario 4 simulations).
@@
 $\bm{\xi}^{x}_{\mathrm{ns}}=\mathbf{0}$.
-For example, $\mathbf{x}$ could be entirely generated from a smooth spatial process, or the dimension $p$ of the spatial effect in (3) could be as large as the sample size $n$ (such as in a regional spatial model with only one data measurement per region).
+For example, $\mathbf{x}$ could be entirely generated from a smooth spatial process, or the dimension $p$ of the spatial effect in (3) could be as large as the sample size $n$ (such as in a regional spatial model with one data measurement per region).
 The covariate vector is then contained in the space spanned by the spatial basis vectors and model (3) is unidentifiable.
@@
 Thus, under this assumption, the frequency components of $\mathbf{x}$ above the threshold can be used to obtain identifiability.
+We note that the frequencies used to obtain identifiability need not be the highest spatial frequencies.
+An example of this is the scenario of Bolin and Wallin, (2024) who consider the asymptotic behaviour of the covariate effect estimate in an analysis model of the form (3) when the true data generation process is given by
+$\mathbf{y}=\beta\mathcal{S}\mathbf{x}+\bm{\epsilon},\quad\bm{\epsilon}\sim N(\bm{0},\sigma^{2}\mathbf{I})$
+where $\mathcal{S}$ is a smoothing operator, for example, $\mathcal{S}\mathbf{x}$ could be the spatial frequencies of the covariate below a threshold $k$ and $\mathbf{x}-\mathcal{S}\mathbf{x}$ the remaining (higher) frequencies.
+Here, there is no spatial confounder as such (in the form of a missing or unmeasured spatial variable), but bias arises due to mis-specification of the covariate in the analysis model.
+However, since $\beta\mathcal{S}\mathbf{x}=\beta\mathbf{x}-\beta(\mathbf{x}-\mathcal{S}\mathbf{x})$, the true data generation process can be written in the form of (2) with $\mathbf{z}=-\beta(\mathbf{x}-\mathcal{S}\mathbf{x})$, and we can therefore use our framework to assess the bias in this case.
+From the expression for $\mathbf{z}$ we see that “confounding” is at the higher frequencies $\mathbf{x}-\mathcal{S}\mathbf{x}$ of the covariate, and the lower frequencies $\mathcal{S}\mathbf{x}$ of the covariate are unconfounded.
+Thus, by Corollary 1, the bias is given by $\frac{-\beta\sum_{i=k+1}^{n}(\xi^{x}_{\textrm{sp},i})^{2}w_{i}}{\sum_{i=1}^{n}(\xi^{x}_{\textrm{sp},i})^{2}w_{i}}$.
+As the sample size $n$ goes to infinity, we expect the higher frequencies to increasingly dominate $\mathbf{x}$ and therefore the bias to approach $-\beta$, consistent with the results of Bolin and Wallin, (2024).
 5.3 Capped spatial+
@@
 As previously noted, unconfounded high frequencies in $\mathbf{x}$ have a similar effect on the bias as non-spatial information.
-Therefore, assuming the $k$ highest frequencies are unconfounded, a natural way to adapt spatial+ is to (i) decompose the covariate as $\mathbf{x}=\mathbf{x}_{\mathrm{sp}}^{k}+\mathbf{r}^{k}$ where $\mathbf{x}_{\mathrm{sp}}^{k}=\sum_{i=1}^{n-k}\xi^{x}_{\mathrm{sp},i}\mathbf{u}_{\mathrm{sp},i}$ and $\mathbf{r}^{k}=\sum_{i=n-k+1}^{n}\xi^{x}_{\mathrm{sp},i}\mathbf{u}_{\mathrm{sp},i}$ with $\mathbf{u}_{\mathrm{sp},i}$ the $i$th spatial eigenvector, and (ii) replace $\mathbf{x}$ by the unconfounded component $\mathbf{r}^{k}$ in model (3).
-Proposition 1 shows that the bias of the corresponding covariate effect estimate is $\langle\mathbf{r}^{k},\mathbf{z}\rangle_{\bm{\mathbf{\Sigma}}^{-1}}/\langle\mathbf{r}^{k},\mathbf{r}^{k}\rangle_{\bm{\mathbf{\Sigma}}^{-1}}$ which should therefore, in theory, be 00 due to the assumption of unconfoundedness.
-However, as the component $\mathbf{r}^{k}$ can also be modelled as part of the highest frequencies of the spatial effect in (3), the estimate of $\beta$ obtained in this way is highly sensitive to inaccuracies at these frequencies of the estimated spatial effect.
+Therefore, assuming the $k$ highest frequencies are unconfounded, a natural way to adapt spatial+ is the following two-step procedure.
+Let $\mathbf{u}_{\mathrm{sp},1},\ldots,\mathbf{u}_{\mathrm{sp},n}$ denote the spatial eigenvectors in $\mathbb{R}^{n}$.
+Decompose the covariate as $\mathbf{x}=\mathbf{x}_{\mathrm{sp}}^{k}+\mathbf{r}^{k}$ where $\mathbf{x}_{\mathrm{sp}}^{k}=\sum_{i=1}^{n-k}\xi^{x}_{\mathrm{sp},i}\mathbf{u}_{\mathrm{sp},i}$ (the lowest $n-k$ frequencies) and $\mathbf{r}^{k}=\sum_{i=n-k+1}^{n}\xi^{x}_{\mathrm{sp},i}\mathbf{u}_{\mathrm{sp},i}$ (the highest $k$ frequencies).
+Estimate the covariate effect $\beta$ by fitting the model (3) but with $\mathbf{x}$ replaced by $\mathbf{r}^{k}$.
+Proposition 1 shows that the bias of the corresponding covariate effect estimate is $\langle\mathbf{r}^{k},\mathbf{z}\rangle_{\bm{\mathbf{\Sigma}}^{-1}}/\langle\mathbf{r}^{k},\mathbf{r}^{k}\rangle_{\bm{\mathbf{\Sigma}}^{-1}}$, which, as $\mathbf{r}^{k}$ is assumed to be unconfounded, should therefore, in theory, be 00.
+However, as the spatial eigenvectors span the whole of $\mathbb{R}^{n}$, $\mathbf{r}^{k}$ and $\mathbf{z}$ no longer lie in orthogonal subspaces.
+As the component $\mathbf{r}^{k}$ can also be modelled as part of the highest frequencies of the spatial effect in (3), the estimate of $\beta$ obtained in this way is highly sensitive to inaccuracies at these frequencies of the estimated spatial effect.
 Such inaccuracies easily arise in practice as the estimation is based on overall spatial smoothing, driven by the fitted values rather than the behaviour at individual spatial frequencies.
@@
 Thus, if we observe stable capped spatial+ estimates for caps below a certain value, these estimates can be expected to be unbiased.
-Finally, we note that the assumption of unconfoundedness does not have to be limited to the highest frequencies of the spatial effect.
+Finally, in Section 5.2 we noted that the assumption of unconfoundedness does not have to be limited to the highest frequencies of the spatial effect.
 If we believe the unconfounded behaviour of $\mathbf{x}$ is in a different frequency range, we can easily adapt the capped spatial+ method by replacing the component $\mathbf{r}^{k}$ by the part of $\mathbf{x}$ that is believed to be unconfounded.
+In the scenario of Bolin and Wallin, (2024) described at the end of Section 5.2, we could use the $k$ lowest frequencies of $\mathbf{x}$ as the component $\mathbf{r}^{k}$ under the assumption that these are contained in the unconfounded low frequency component $\mathcal{S}\mathbf{x}$ of the covariate.
+This is therefore also consistent with the suggestion in Bolin and Wallin, (2024) to replace $\mathbf{x}$ in the analysis model (3) by a smoothed (i.e.
+lower frequency) version $\hat{\mathcal{S}}\mathbf{x}$ of the covariate.
 5.4 Simulations for capped spatial+
@@
 $\mathbf{y}=\beta\mathbf{x}+\mathbf{z}+\bm{\epsilon}^{y},\quad\mathbf{x}=\mathbf{z}^{x}_{\mathrm{sp,low}}+\mathbf{z}^{x}_{\mathrm{sp,medium}}+a\ \mathbf{z}^{x}_{\mathrm{sp,high}},\quad\mathbf{z}=\mathbf{z}^{z}_{\mathrm{sp,low}}+\mathbf{z}^{x}_{\mathrm{sp,medium}}$
-where $\bm{\epsilon}^{y}\sim N(\mathbf{0},\sigma^{2}\mathbf{I})$ and $a$ measures the weight of the unconfounded high frequencies in $\mathbf{x}$.
+where $\bm{\epsilon}^{y}\sim N(\mathbf{0},\sigma^{2}\mathbf{I})$ and $a>0$ controls the size of the unconfounded high frequency spatial component in $\mathbf{x}$.
 In order to generate $\mathbf{z}^{x}_{\mathrm{sp,low}}$, $\mathbf{z}^{x}_{\mathrm{sp,high}}$ and $\mathbf{z}^{x}_{\mathrm{sp,medium}}$, we fit a thin plate spline to a realization of a Gaussian process with $\kappa=0.5/3$ (see beginning of the section), where the penalty and design matrix have now been reparameterised such that they only consider the $100$ lowest, the $k$ highest and all medium frequencies, respectively; $\mathbf{z}^{z}_{\mathrm{sp,low}}$ is generated similarly to Section 4 but it is fitted to a new realization of a Gaussian process with $\kappa=0.5/3$ using only the $100$ lowest basis functions.
 We consider $k\in\{5,10,20\}$ and $a\in\{1,2\}$ in the data generating process; $\beta=1$, $\sigma=0.1$, $n=1000$, $N=50$.
-The reparametrisation is explained in Web Appendix 2.
+The reparametrisation is explained in Appendix 2.
 In the data analysis model, the capped spatial+ model uses a reparameterised $\tilde{\mathbf{B}}_{\mathrm{sp}}$ and $\tilde{\bm{\beta}}_{\mathrm{sp}}$ in the design and penalty matrix for the spatial effects in the first and second stage equation of spatial+ with $\mathrm{cap}\in\{5,10,15,20,25,30,40,50\}$, so that we consider cases where the used cap is above, below or equal to the true cap $k$.
 Unlike the standard spatial+, this allows us to control the frequencies included in the spatial effects for $\mathbf{y}$ and $\mathbf{x}$, such that ideally $\mathbf{r}^{x}$ should not have any confounded frequencies.
-Crucially here, on the first stage equation for $\mathbf{x}$ we use no smoothing in order to guarantee the residuals $\mathbf{r}^{x}$ are orthogonal to the subspace spanned by $\tilde{\mathbf{B}}_{\mathrm{sp}}\tilde{\bm{\beta}}_{\mathrm{sp}}$.
+Here, on the first stage regression for $\mathbf{x}$ we use no smoothing in order to guarantee the residuals $\mathbf{r}^{x}$ are orthogonal to the subspace spanned by $\tilde{\mathbf{B}}_{\mathrm{sp}}\tilde{\bm{\beta}}_{\mathrm{sp}}$.
 We use $n$ basis functions in the spatial model.
-In Figure 3, the bias decreases for increasing $k$ and $a$, as larger $k$ implies less confounded medium frequencies and more unconfounded high frequencies, and larger $a$ implies more weight being put on unconfounded high frequencies in $\mathbf{x}$ and thus less bias.
+In Figure 3 we see that the bias decreases for increasing $k$ and $a$.
+This is expected from our results as a larger $k$ implies fewer confounded medium frequencies and more unconfounded high frequencies, while a larger $a$ implies a larger proportion of unconfounded high frequencies in $\mathbf{x}$.
 For $\mathrm{cap}\leq k$, we are able to completely remove confounding bias, even if only using a smaller fraction of the unconfounded high frequency component in $\mathbf{x}$ to identify the effect.
@@
 Thus, we observe a constant bias of zero for caps below $k$ and a considerable increase for caps above $k$.
-The MSE for the predicted values reveals that in general capping increases MSE, although quite negligibly for $k\leq\mathrm{cap}$, when compared to the drastic increase observed for $k>\mathrm{cap}$ (see Web Appendix 3.1).
-Figure 3: Bias of $\widehat{\beta}$ in spatial model and capped spatial+ models with different combinations of $k\in\{5,10,20\}$ and $a\in\{1,2\}$.
-The true value is $\widehat{\beta}=1$.
+The MSE for the predicted values reveals that in general capping increases MSE, although quite negligibly for $k\leq\mathrm{cap}$, when compared to the drastic increase observed for $k>\mathrm{cap}$ (see Appendix 3.1).
+Figure 3: Bias of $\widehat{\beta}$ in the spatial and capped spatial+ models with different combinations of $k\in\{5,10,20\}$ and $a\in\{1,2\}$.
+The true value is $\beta=1$.
 All other parameters remain constant.
 6 Application to air temperature in Germany
-As a motivating study, we analyse the relationship between the monthly mean temperature at two metres above ground, treated here as the response, and the monthly mean precipitation, the covariate, in Germany for the year 2010 at 336 locations.
+As an empirical illustration, we analyse the relationship between the monthly mean temperature at two metres above ground, treated here as the response, and the monthly mean precipitation, the covariate, in Germany for the year 2010 at 336 locations.
 The data are open access and provided by the German Meteorological Service (DWD).
@@
 By comparing the results of (1) and (2), we show how capped spatial+ can help identify problematic covariate effect estimates in the spatial model and suggest appropriate adjustments.
-All spatial effects are modelled with thin plate spline based on mgcv.
+All spatial effects are modelled with thin plate regression splines in mgcv (Wood,, 2024).
 The results are shown in Figure 4.
-As the cap used in our spatial+ increases, the fit of the model deteriorates, so we only show the estimates with $cap=5,6,\ldots,15$ (caps below 5 may not be reliable as there is likely little spatial information left in the residuals), which are also statistically significant at the 95% confidence level.
+When the cap used for capped spatial+ is very small, estimates of the covariate effect may not be reliable, as the residuals will be very small and unlikely to have sufficient information about the covariate.
+On the other hand, as the cap increases, the fit of the model deteriorates.
+Therefore, we only show the estimates with $cap=5,6,\ldots,15$ which are also statistically significant at the 95% confidence level.
 From September to January and May, the caps (below 15) are generally insignificant, as indicated by the lack of pink dots in the plot.
-In the other months, the capped spatial+ can be used to assess whether we have confidence in the spatial estimate, and the dots show how the estimates should likely be adjusted.
-November is perhaps the most puzzling, where the coefficients appear to be either positive or negative depending on the frequencies included in the capped spatial+.
+In the other months, the capped spatial+ estimates can be used to assess whether we have confidence in the spatial estimate, and the dots show how the estimate should likely be adjusted.
+November is perhaps the most puzzling, where the coefficients appear to be either positive or negative depending on the frequencies included in capped spatial+.
 On closer inspection, this can be explained by a more complex spatial structure of precipitation in this month, with many localised patterns, so that a linear effect of precipitation may not be sufficient in general.
@@
 In order to understand the observed behaviour in more detail, we focus on three months: January, April and August.
-The results are shown in Table 1 for caps up to 9, which also shows the AIC and RMSE of the different models.
+For these three months, results for caps up to 9, including the AIC and RMSE of the different models, are shown in Table 1.
 We start with the spring and summer months.
@@
 This suggests that unobserved spatial confounders are not affecting the estimation and the spatial model estimate does not need to be adjusted.
-In January, we know from Figure  that none of the capped spatial+ estimates up to a cap of 15 are statistically significant.
+In January, we know from Figure 4 that none of the capped spatial+ estimates up to a cap of 15 are statistically significant.
 This suggests that the highest frequencies (up to 15) in the spatial effect are not helping us to identify the effect of rainfall and may simply not be present in this particular covariate.
@@
 In practical applications, the difference between the estimates in the spatial and non-spatial models is commonly seen as an indicator of spatial confounding.
-However, our analysis and simulations show that this difference is not directly related to the size or occurrence of bias in the spatial model.
+However, our analysis and simulations show that a difference between these estimates does not in itself show whether the estimate in the spatial model is biased.
 Instead we propose using the methods set out in Section 5 to assess and adjust for the bias.
 As most other work on spatial confounding, the focus in this paper has been on quantifying the bias in spatial regressions.
-A derivation and short discussion about the variance of the covariate effect estimate is included in Web Appendix 1.6.
+A derivation and short discussion about the variance of the covariate effect estimate is included in Appendix 1.6.
 However, to put the bias into perspective and thoroughly conduct statistical inference, more investigations on associated uncertainty will be an important direction for future research.
 Data availability statement
-The data is publicly available on the website of the German weather service (DWD).
-Supporting Information
-Web Appendices referenced in Section 26, are included below.
-See pages - of appendix.pdf
+The data is publicly available on the website of the German weather service (DWD). We provide code for all analyses in the paper, which includes instructions on how to download the data.